A Bibliography of Ada User and Ada User Journal

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

06 January 2024 Version 1.18

Title word cross-reference

$5.6 [425].

11 [289]. 12th [337]. 13th [455]. 14th [618, 617]. 15th [728, 660]. 162 [121]. 16th [744]. 178B [708]. 17th [875]. 1998 [231, 257].

2 [41]. 2001 [250, 260, 268, 276]. 2002 [300, 309, 282, 293, 317, 324]. 2003 [331]. 2004 [346, 356, 365, 373, 380]. 2005 [389, 399, 409, 417, 330, 372, 386, 406, 396, 379, 405, 387, 395, 415, 487, 496, 525, 306, 536, 628, 488, 499, 497, 500, 486, 474]. 2006 [426, 434, 442, 449, 424]. 2007 [457]. 2009 [544]. 200Y [323, 585]. 2010 [609, 622, 632]. 2011 [641, 663]. 2012 [826, 686, 700, 679, 717, 873, 658, 647, 669, 707, 680, 693, 727, 742, 718, 874, 659]. 2013 [736, 748, 764, 777, 733]. 2014 [811, 785, 837, 774, 802, 761, 838]. 2015 [860]. 21st [176]. 24772 [535].

3 [7, 46].

6a [406, 727]. 6th [20].

8-bit [710]. 83 [96, 638, 292, 162, 91, 104]. 8652 [344, 343, 342].

95 [134, 168, 157, 156, 158, 122, 197, 316, 292, 174, 456, 299, 239, 307]. 9X [119, 59, 78, 139, 96, 97, 98, 102, 65, 159, 91, 104, 67].

AADL [572]. abstract [90]. Abstraction [274, 584, 17]. Abstractions [878]. academia [235]. Academic [191, 353, 140]. Access [156, 386]. achieves [131]. Achieving [89]. activities [23]. Ada [20, 129, 605, 631, 640, 651, 662, 627, 672, 685, 699, 735, 747, 763, 762, 751, 728, 793, 785, 875, 231, 618, 744, 337, 61, 69, 617, 660, 569, 207, 455, 142, 117, 80, 649, 826, 59, 78, 20, 26, 30, 42, 54, 93, 103, 135, 139, 151, 140, 133, 134, 131, 163, 177, 166, 180, 168, 188, 189, 190, 198, 199, 208, 191, 200, 193, 202, 213, 214, 237, 215, 238, 217, 236, 219, 221, 222, 249, 259, 267, 275, 250, 260, 268, 276, 241, 251, 261]. Ada [269, 300, 309, 282, 293, 277, 283, 294, 301, 317, 324, 331, 310, 318, 325, 332, 340, 346, 356, 365, 373, 347, 357, 366, 341, 380, 389, 399, 409, 374, 381, 390, 400, 417, 426, 434, 442, 410, 418, 427, 435, 466, 478, 449, 457, 477, 443, 450, 458, 467, 490, 503, 518, 528, 489, 502, 517, 527, 479, 491, 504, 519, 538, 550, 563, 574, 537, 549, 562, 529, 539, 551, 564, 587, 599, 608, 621, 609, 622, 598, 607, 620, 593, 575, 588]. Ada [600, 610, 594, 631, 640, 651, 662, 632, 641, 652, 663, 630, 639, 661, 623, 633, 642, 653, 711, 672, 685, 699, 673, 686, 700, 671, 684, 698, 664, 674, 687, 679, 722, 735, 747, 763, 736, 748, 764, 721, 734, 746, 717, 755, 818, 776, 792, 803, 819, 777, 804, 775, 783, 811, 828, 839, 850, 866, 829, 840, 851, 867, 860, 880, 881, 873, 353, 2, 524, 596, 64, 96, 97, 98, 102, 115, 157, 156, 158, 230, 122, 123, 330, 372, 386, 406, 396, 379, 405]. Ada [387, 395, 415, 534, 658, 647, 669, 707, 680, 693, 727, 742, 718, 858, 874, 73, 141, 197, 29, 85, 650, 425, 516, 23, 638, 184, 185, 1, 235, 371, 509, 119, 19, 28, 65, 266, 316, 487, 496, 606, 619, 15, 145, 8, 825, 62, 292, 525, 70, 32, 162, 53, 697, 7, 174, 424, 544, 92, 4, 41, 5, 75, 256, 456, 94, 47, 659, 14, 113, 45, 68, 72, 107, 246, 6, 40, 159, 299, 24, 114, 13, 105, 35, 306, 22, 27, 31, 39, 50, 89, 66, 87, 81, 90, 440, 76, 397]. Ada [10, 38, 338, 146, 514, 16, 173, 510, 558, 91, 352, 362, 536, 628, 84, 239, 546, 290, 147, 801, 9, 95, 488, 33, 143, 388, 416, 499, 512, 710, 696, 857, 573, 720, 733, 363, 515, 719, 291, 307, 770, 86, 211, 110, 18, 175, 570, 597, 398, 25, 323, 585, 104, 556, 63, 497, 759, 148, 501, 55, 67, 500, 486, 227, 43, 57, 58, 132, 226, 160, 474, 692, 701, 712, 723, 737, 749, 765, 778, 794, 805, 820, 830, 841, 852, 868]. Ada/C [597]. Ada9X [79]. Adaland [756]. Adaptability [247]. AdaUK [178]. AdaWay [706]. Additional [330]. AdDoc [810]. Adjusting [197]. Advanced [720]. affiliates [140]. Agenda [340]. Aircraft [849]. Alan [258]. Algebraic [117, 90, 95]. algebras [73]. Algorithm [787, 307, 162]. Alice [756]. Alignment [770]. Allocating [500]. Alternative [173]. Amendment [344, 343, 342]. Analysis [464, 547, 571, 849, 484, 847, 836, 345, 99, 86]. Analytical [846]. Andy [258]. Annex [619]. Annual [593]. APIs [288]. applets [227]. Application [731, 791, 484, 147]. Applications [571, 546, 827, 572, 92, 105, 84, 209]. Approach [786, 307, 513, 117, 7]. Approaches [811, 860, 816]. Approved [717, 873, 69]. Architecture [849]. Architecture-led [849]. Architectures [670, 144]. ARG [343, 342]. Argument [559]. Arguments [859]. Ari [569]. Arrays [257]. ASIM [681]. ASIS [208]. ASISFAQ [198]. ASISRG [204]. ASISWG [204]. ASISWG/ASISRG [204]. Asked [198, 208]. aspects [658]. Asphodel [45]. Assembly [340, 341, 230]. ASSERT [709]. Associate [249, 259, 267, 275, 466, 478, 490, 503, 518, 528, 538, 550, 563, 574, 587, 599, 608, 621, 631, 640, 651, 662, 711, 672, 685, 699, 722, 735, 747, 763, 818, 776, 792, 803, 828, 839, 850, 866, 880]. Assurance [813, 848, 682]. Asynchronous [65]. ATAC [109]. automatic [101]. Automation [848]. Avionics [465, 638, 66]. AVR [801, 710]. AVR-Ada [801]. Award [692]. Awareness [849]. AWS [363].

Babbage [835]. backpropagation [162]. Bare [497]. Based [816, 545, 488, 827, 501, 117, 7, 859, 107, 81, 836, 572]. Basic [398]. be [185]. Beat [706, 692]. Before [345]. beginner [227]. beginners [118]. Behavioral [815]. Behaviour [786]. being [22]. Ben [569]. Ben-Ari [569]. Benchmarking [448]. Benefits [682, 408]. between [96, 110]. Beyond [176, 499, 500, 810]. Bias [557]. bibliography [93]. Billion [425]. binding [1]. Bindings [597]. Bioinformatics [770]. bit [710]. Board [189, 497, 773]. Book [164, 258, 569]. Branch [398]. Breaks [345]. brief [51]. Briefing [188]. Bring [440]. Budgets [498]. Buffers [536]. Bug [397]. Building [546, 476]. builds [407]. Burns [258]. business [472]. buyer [54]. Byron [858].

C [524, 68, 72, 110, 597, 160]. CAIS [11, 12]. Calendar [232, 244, 254, 264, 272, 280, 286, 297, 304, 313, 321, 328, 335, 350, 360, 369, 377, 384, 393, 403, 413, 421, 430, 438, 446, 453, 461, 470, 482, 494, 507, 522, 532, 542, 554, 567, 578, 591, 603, 613, 626, 636, 645, 656, 704, 667, 677, 690, 715, 740, 753, 808, 768, 781, 797, 833, 844, 871, 823, 855]. Call [288]. Callbacks [597]. Can [123]. capture [161]. Case [465, 791, 848, 847, 784, 66, 110, 63]. century [176]. Certification [463]. Chairman [169, 221]. Challenge [706, 692]. Challenges [811, 860, 862]. Change [789]. changes [134]. Channel [100]. Charles [364]. charts [29]. Checking [719]. checkpointing [209]. choice [68, 72, 105]. Classes [800]. client [70]. club [138]. Co [709]. Co-design [709]. Code [197, 697, 299, 836, 463, 838, 681, 488, 101]. codebase [743]. ’Coding [580, 526, 719]. collaborative [19]. colleges [94]. Combining [697]. Command [525, 46]. Comments [343]. commercial [108, 147]. committee [133, 129, 171, 183, 191]. communication [149, 212]. Community [594]. Comparing [160]. Comparison [526, 68, 72, 628]. Compatibility [96]. competition [213]. Compilation [119, 743]. Compiler [197, 463]. Compilers [26, 30, 37, 54, 104]. complex [638, 472]. Component [827, 572, 6]. Component-Based [827, 572]. components [7, 5, 45, 90, 416, 63]. Computation [307]. computer [21, 18]. Computing [864, 879, 825]. CONCERTO [865]. Conclusions [617, 511]. Concurrency [560, 827, 732, 106, 161]. Conference [128, 232, 244, 254, 264, 272, 280, 286, 297, 304, 313, 321, 328, 335, 350, 360, 369, 377, 384, 393, 403, 413, 421, 430, 438, 446, 453, 461, 470, 482, 494, 507, 522, 532, 542, 554, 567, 578, 591, 603, 613, 626, 636, 645, 656, 704, 667, 677, 690, 715, 740, 753, 808, 768, 781, 797, 833, 844, 871, 823, 855, 424, 544, 338, 733, 20, 103, 142, 231]. Conformance [876]. conformant [121]. Cons [561]. Considered [513, 580, 77]. construction [81]. Consumption [815]. Container [364]. Containers [826, 406, 727, 774]. Contest [593, 605, 627]. Context [596, 859]. continuing [87]. Contract [816]. Contract-Based [816]. Contracts [815, 658]. Contribution [256]. Control [525, 659, 536, 398, 65, 46, 730]. Controlled [813]. Controllers [710]. conversion [41]. converting [91]. Coq [770]. CORBA [440, 209]. Correcting [474]. Correctness [274, 584]. Corrigendum [873]. cosmology [27]. Cost [42, 116]. Costs [545, 160]. COTS [308]. countess [41]. Counting [116]. Couverture [571]. Coverage [571]. CPS [813, 816, 865]. CPU [82]. Creating [132]. Creation [835]. Crimeville [649]. Critical [354, 596, 571, 799]. Criticality [787, 785, 791, 790, 814, 786, 789, 788]. Cross [546]. CubeSat [516]. current [59]. curriculum [18]. Customizing [299]. Cyber [811, 860, 864, 879, 863, 862]. Cyber-Physical [811, 860, 864, 879, 863, 862].

DAMOKLES [52]. Data [802, 81, 648, 34, 89, 90, 17]. database [52]. Days [172]. DDC [85]. DDC-I [85]. Deadlock [239]. Deal [816]. debug [113]. debugging [111]. Defence [87, 23]. defense [92]. Defining [344]. Depend [123]. Dependable [811, 860, 106, 814, 863]. Dependencies [802]. Deploying [865]. Deriving [859]. Design [174, 46, 456, 3, 80, 119, 709, 99, 114, 66, 17]. Designer [432, 586]. Designing [865, 719, 670, 629]. designs [6]. Detector [239]. Develop [226]. Developing [433, 363, 16]. Development [256, 299, 84, 398, 160, 773, 100, 29, 1, 145, 176, 14, 696, 74]. developments [11, 408]. device [595]. Diagrams [784]. difficult [38]. Digest [822, 854, 577, 590, 602, 612, 625, 635, 644, 655, 703, 666, 676, 689, 714, 725, 739, 752, 807, 767, 780, 796, 832, 843, 870]. Distributed [113, 246, 299, 536, 476, 149, 144]. Distribution [173, 247, 616]. Do [257, 708]. DO-178B [708]. DO178C [847]. Document [557, 810]. Documented [784]. DoD [185]. does [234]. don’t [186, 582, 39]. download [472]. Driven [709, 456, 681]. drivers [595]. dual [120].

early [159]. eBook [783]. ECMA [121]. ECMA-162 [121]. EDF [474]. Editor [195]. Editorial [124, 136, 152, 165, 179, 166, 180, 193, 202, 192, 201, 216, 217, 236, 241, 251, 261, 269, 242, 252, 262, 270, 277, 283, 294, 301, 278, 284, 295, 302, 310, 318, 325, 332, 311, 319, 326, 333, 347, 357, 366, 348, 358, 367, 374, 381, 390, 400, 375, 382, 391, 401, 410, 418, 427, 435, 411, 419, 428, 436, 443, 450, 458, 467, 444, 451, 459, 468, 479, 491, 504, 519, 480, 492, 505, 520, 529, 539, 551, 564, 530, 540, 552, 565, 575, 588, 600, 610]. Editorial [576, 589, 601, 611, 623, 633, 642, 653, 624, 634, 643, 654, 702, 664, 674, 687, 701, 665, 675, 688, 712, 723, 737, 749, 713, 724, 738, 750, 806, 765, 778, 794, 805, 766, 779, 795, 820, 830, 841, 852, 831, 842, 869, 868, 228, 812, 861, 821, 853]. Editors [33]. Education [514, 512, 515, 513, 17, 211]. effect [104]. Effective [847]. Efficiency [814]. efficient [159]. Eight [207]. Embedded [484, 74]. emption [22]. Energy [815]. enforcing [82]. Engine [846]. Engineering [811, 860, 709, 514, 290, 512, 513, 52, 211]. Engineers [569]. Enhancements [323, 585]. Entails [425]. Entity [670]. Entity-Life [670]. Enumerating [561]. Environment [157, 246, 743, 70, 52, 113, 148, 119, 210]. environments [21, 13]. Epilogue [415, 742]. ERAM [524]. ERB [448]. etc [396, 707]. Ethernet [790]. Europe [300, 309, 317, 324, 331, 346, 356, 365, 373, 380, 389, 399, 409, 417, 426, 434, 442, 449, 457, 609, 622, 631, 640, 651, 662, 632, 641, 652, 663, 672, 685, 699, 673, 686, 700, 735, 747, 763, 736, 748, 764, 777, 793, 811, 785, 860, 231, 230, 556, 189, 219, 222, 249, 259, 267, 275, 340, 341, 466, 478, 490, 503, 518, 528, 538, 550, 563, 574, 587, 599, 608, 621, 711, 722, 751, 818, 776, 792, 803, 819, 804, 828, 839, 850, 866]. Europe [829, 840, 851, 867, 880, 881, 338]. European [799]. Evaluating [464, 570, 43]. evaluation [10, 9]. evaluator [44]. Events [153, 167, 181, 194, 203, 233, 245, 255, 265, 273, 281, 287, 298, 305, 314, 322, 329, 336, 351, 361, 370, 378, 385, 394, 404, 414, 422, 431, 439, 447, 454, 462, 471, 483, 495, 508, 523, 533, 543, 555, 568, 579, 592, 604, 614, 637, 646, 657, 705, 668, 678, 691, 716, 726, 741, 754, 809, 769, 782, 798, 824, 834, 845, 856, 872, 127, 218]. Evolutionary [307]. Exceptions [158, 396]. Executable [596]. Executing [827]. Execution [465, 487, 547, 659, 847, 498, 82]. Execution-Time [487]. executive [80]. Experience [496, 595, 440, 773, 162]. Experiences [354, 197, 83, 514, 570]. experiment [159]. experiments [40]. Expressions [813, 669]. Expressive [316]. Extensibility [787]. Extraction [784].

Families [71]. family [212]. Fault [548, 159, 209]. Feasibility [816]. Feature [784]. Features [659, 497]. Field [862]. Final [134]. Finder [397]. Fine [877]. Fine-grained [877]. First [692, 338, 799, 162, 6]. Five [248]. Fixing [345]. Flexibility [814]. Flight [836]. Flow [802]. Formal [774, 45]. Forthcoming [127, 153, 167, 181, 194, 203, 218, 233, 245, 255, 265, 273, 281, 287, 298, 305, 314, 322, 329, 336, 351, 361, 370, 378, 385, 394, 404, 414, 422, 431, 439, 447, 454, 462, 471, 483, 495, 508, 523, 533, 543, 555, 568, 579, 592, 604, 614, 637, 646, 657, 705, 668, 678, 691, 716, 726, 741, 754, 809, 769, 782, 798, 824, 834, 845, 856, 872]. Framework [448, 571, 441, 486, 62, 16]. Frameworks [731]. Free [783, 408]. Frequently [198, 208]. Frog [220]. FTP [118]. fully [121]. Functional [838]. Future [862].

game [649]. GEC [47]. Gems [477, 489, 502, 517, 527, 537, 549, 562, 598, 607, 620, 630, 639, 661, 671, 684, 698, 721, 734, 746, 762, 775]. General [340, 341, 230, 75]. Generate [197]. Generating [416, 572]. Generation [864, 879, 697, 299, 681, 488, 150, 101]. generator [810]. Generic [257]. Generics [697, 683, 396]. Genetic [307]. Ghost [838]. GNAT [130, 743, 771, 159, 315, 463, 581]. grained [877]. graphic [53]. Graphical [1]. Great [355]. Group [36]. Groups [498]. Guide [650, 266, 710, 573, 54, 227]. Guided [315]. Guidelines [5, 388].

Handling [789]. Hardware [709, 3]. Hardware/Software [709]. Harmful [513, 580, 77]. harness [75]. Heap [745]. Help [345]. heterogeneous [149]. HI [572]. hiding [89]. Hierarchical [122, 102]. High [308, 583, 682, 863, 648, 476, 266, 145, 108, 71]. High-Assurance [682]. High-Integrity [476, 145]. High-Risk [863]. highly [68, 72]. history [581]. Home [863]. HRG [266].

Identifying [465]. IEC [257, 56, 344, 343, 535, 342]. IEEE [21]. II [22]. IMAGE [34]. IMAGES [34]. impact [146, 91]. implement [697]. Implementation [496, 659, 10, 801, 629, 648, 28, 121]. Implementing [638, 159, 95, 497, 162]. implications [42, 18]. Important [515]. Impressions [289]. Improvements [759]. improving [416]. Including [709]. incompatibilities [91]. Increasing [316]. incremental [70]. Industrial [785, 526]. Industrials [408]. industry [695]. Information [151, 177, 190, 199, 214, 237, 802, 135, 163]. Infrastructure [20]. Inherent [862]. inherited [743]. Initiative [353]. Innovative [571]. Insecurities [57]. Instructions [343, 342]. integrate [60]. Integrating [246, 710, 475]. Integrity [308, 583, 648, 476, 266, 145, 108]. Interface [198, 208, 12]. interfaces [212, 13, 11]. Interfacing [524, 510]. International [20, 103, 728, 875, 618, 744, 337, 617, 660, 207, 455, 131, 49, 129]. Interworking [110]. Introducing [47]. Introduction [219, 372, 647, 6]. Investment [425]. Invitation [168, 306]. IPSE [51]. IRTAW [728, 875, 660, 455, 289, 511]. IRTAW-13 [455]. IRTAW-15 [660]. IRTAW14 [618]. IRTAW8 [207, 206]. ISO [679, 717, 873, 257, 344, 343, 535]. ISO/IEC [257, 344, 343, 535, 342]. issue [22, 27, 31, 39, 50]. Issues [876, 729, 760, 248, 473, 616, 730, 732, 80, 28, 69, 39]. Iterators [707]. IV [31].

Java [800, 235, 371, 291, 513, 227]. Journal [166, 180, 193, 202, 217, 236, 241, 251, 261, 269, 277, 283, 294, 301, 310, 318, 325, 332, 347, 357, 366, 374, 381, 390, 400, 410, 418, 427, 435, 443, 450, 458, 467, 479, 491, 504, 519, 529, 539, 551, 564, 575, 588, 600, 610, 623, 633, 642, 653, 664, 674, 687, 534, 701, 712, 723, 737, 749, 765, 778, 794, 805, 820, 830, 841, 852, 868]. JSD [15, 35].

Kernel [1, 497]. Kissing [220]. know [186, 582]. Kron [814]. Kron-OS [814].

lady [38]. lang [788]. Language [594, 679, 717, 755, 783, 873, 858, 731, 561, 256, 432, 586, 558, 535, 546, 557, 473, 559, 616, 878, 131, 68, 72, 105, 234, 86, 694, 629]. Languages [3, 247, 186, 71, 258]. Large [292, 743, 38]. law [22, 176]. Layer [212]. Leadership [155]. Leaks [800]. Least [425]. led [849]. Legacy [440, 148]. LEGO [650]. Lessons [155]. Letter [594, 751]. Letters [195]. Level [708, 847]. Liaison [129]. Libraries [485]. Library [122, 718, 102, 405, 364, 226]. Life [440, 670]. Line [638, 697, 649]. Linear [789]. Lines [398, 111]. list [37]. Living [407]. LNCS [783]. local [19]. Locking [758]. LOTOS [76]. Lovelace [825, 835, 857]. Low [58]. Low-Ada [58]. Lower [212]. Ltd [47].

maintain [743]. Maintenance [594, 545]. Major [248]. Making [209]. MALPAS [55]. Management [799]. managing [17]. mandated [185]. Mapping [15, 114]. Market [425, 256]. marketing [133, 171, 183]. MaRTE [573]. MASCOT [46, 7]. matter [234]. mature [104]. Me [706, 692]. means [60]. Measuring [82]. mechanism [159]. Medical [863]. Meet [187]. Meeting [204]. meets [41]. Members [249, 259, 267, 275, 466, 478, 490, 503, 518, 528, 538, 550, 563, 574, 587, 599, 608, 621, 631, 640, 651, 662, 711, 672, 685, 699, 722, 735, 747, 763, 818, 776, 792, 803, 828, 839, 850, 866, 880]. Memories [432, 586]. Message [169, 221, 222]. Metamodel [488]. Metamodel-Based [488]. Methodologies [862]. Methodology [865]. methods [89]. metrics [138, 74]. Micro [710]. Micro-Controllers [710]. military [88]. MINDSTORMS [650]. missing [475]. Mixed [787, 785, 791, 790, 814, 786, 788]. Mixed-Criticality [787, 786]. MOD [10, 248]. Mode [789]. Model [709, 848, 784, 681, 379, 83]. Modeling [862, 670]. modelling [2]. Models [859]. Modula [41]. Modula-2 [41]. Monitoring [536, 82]. Mordechai [569]. Motions [341]. multi [649, 407]. multi-user [649]. Multicomputer [173]. Multicore [757]. Multiprocessor [615, 729, 730]. Murphy [22]. My [162, 338]. MyCCM [572]. MyCCM-HI [572].

National [249, 259, 267, 275, 466, 478, 490, 503, 518, 528, 538, 550, 563, 574, 587, 599, 608, 621, 631, 640, 651, 662, 711, 672, 685, 699, 722, 735, 747, 763, 818, 776, 792, 803, 828, 839, 850, 866, 880]. NATO [25]. Naval [525]. Need [848]. nets [32]. network [2, 15]. Networks [790]. News [822, 854, 126, 137, 154, 171, 170, 182, 196, 205, 223, 229, 243, 253, 263, 271, 279, 285, 296, 303, 312, 320, 327, 334, 339, 349, 359, 368, 376, 383, 392, 402, 412, 420, 429, 437, 445, 452, 460, 469, 481, 493, 506, 521, 531, 541, 553, 566, 577, 590, 602, 612, 625, 635, 644, 655, 703, 666, 676, 689, 714, 725, 739, 752, 807, 767, 780, 796, 832, 843, 870]. Next [864, 879, 352, 248, 511]. Next-Generation [864, 879]. No [355]. node [107]. non [92]. non-defense [92]. Note [56]. ’Notes [835, 846, 125]. nothing [186, 582]. Numeric [330]. numerical [73]. NXT [650].

O2 [141]. Object [97, 379, 174, 99, 463, 838, 247, 323, 585, 226, 62, 88, 114]. Object-Oriented [463, 247, 323, 585, 226, 97, 99, 62, 88, 114]. objects [98, 144]. Ocarina [572]. off [108]. off-the-shelf [108]. old [73]. old-numerical [73]. On-board [773]. on-line [649]. One [514]. OO [175, 423]. OO-programming [175]. OODBMS [141]. OOP [475]. Open [571, 865, 760]. Opening [64]. Operating [865, 510, 498, 21]. Opinion [580]. Opportunities [465]. Optimization [547]. order [71]. Organisations [249]. Organizations [259, 267, 275, 466, 478, 490, 503, 518, 528, 538, 550, 563, 574, 587, 599, 608, 621, 631, 640, 651, 662, 711, 672, 685, 699, 722, 735, 747, 763, 818, 776, 792, 803, 828, 839, 850, 866, 880]. Oriented [174, 463, 838, 247, 323, 585, 226, 97, 379, 62, 99, 88, 114]. Origins [581]. Other [509]. our [694]. Overview [728, 875, 12, 17, 51]. Oxford [857].

package [2]. Packages [330, 34, 14, 143]. Packaging [101, 89]. page [179]. Panel [745]. Papers [224]. Paraffin [628]. Parallel [757, 501, 629]. Parallelism [4, 877, 628]. ParaSail [629]. Paris [353]. part [472, 729, 730]. Partitioned [787]. Partitioning [787]. partitions [83]. passive [79]. Pattern [545, 423]. Pattern-Based [545]. Patterns [174, 485, 488]. PCTE [49, 11, 12, 48]. pedagogy [176]. persistence [141]. Persistent [826]. perspective [695, 24]. Physical [811, 860, 864, 879, 771, 863, 862]. Pilot [570]. plan [78]. Planning [88]. Plans [511]. Platform [836]. Platforms [864, 879]. Pointers [355, 240]. policies [101]. Policy [166, 180, 193, 202, 217, 236, 241, 251, 261, 269, 277, 283, 294, 301, 310, 318, 325, 332, 347, 357, 366, 374, 381, 390, 400, 410, 418, 427, 435, 443, 450, 458, 467, 479, 491, 504, 519, 529, 539, 551, 564, 575, 588, 600, 610, 623, 633, 642, 653, 664, 674, 687, 701, 712, 723, 737, 749, 765, 778, 794, 805, 820, 830, 841, 852, 868, 185]. Polymorphic [220, 597]. Pools [707]. portable [21]. Porting [525]. POSIX [21, 149]. Postbank [147]. Power [316]. Practical [354, 484]. Practicum [155]. pre [22]. pre-emption [22]. Predefined [157, 405, 718, 77]. Predictable [66]. Pregnancy [863]. Preparation [343, 342]. President [222, 751]. Press [679, 692, 755, 783, 873]. Principled [513]. Privacy [800]. Probabilistic [786]. problem [161]. Proceedings [811, 785, 860]. Process [155, 709, 859, 472]. Process-based [859]. Processors [500]. Procurement [248]. Product [638, 697]. professional [19]. Professor [514]. Profile [731, 29, 595, 266, 583]. Program [173, 14]. programme [19]. Programmers [371]. Programming [213, 593, 605, 594, 627, 487, 561, 485, 789, 558, 535, 838, 720, 258, 557, 291, 323, 585, 97, 186, 62, 68, 72, 595, 175]. Programs [355, 583, 239, 27, 81]. Progress [618, 535]. Project [59, 516, 292, 570, 78, 407, 51, 606, 619, 865, 681]. projects [99, 88, 91]. Proof [683]. Properties [816]. Proposal [344]. Proposed [341]. Pros [561]. Protected [98, 79]. protocol [474]. Protocols [758, 247, 212, 730]. prototyping [32, 112]. PROXIMA [786]. Published [755, 783]. Publisher [441]. purpose [75].

Quagmire [345]. Qualification [772]. Quality [155, 836, 345]. Quarterly [822, 854, 577, 590, 602, 612, 625, 635, 644, 655, 703, 666, 676, 689, 714, 725, 739, 752, 807, 767, 780, 796, 832, 843, 870]. Questions [198, 208]. queueing [2, 86]. queuing [132].

Rail [799]. Railway [791, 859]. Rapid [32, 112]. Rapidly [433]. rather [235]. Rational [210]. Rationale [372, 386, 406, 396, 379, 405, 387, 395, 415, 658, 647, 669, 707, 680, 693, 727, 742, 718, 774, 802, 761, 838, 71, 755]. Ravenscar [266, 837, 448, 583, 595, 476]. re [6]. re-use [6]. Reactive [299, 670]. Real [728, 875, 395, 693, 371, 316, 618, 744, 814, 337, 583, 745, 660, 720, 258, 291, 497, 67, 207, 486, 455, 80, 149, 7, 99, 113, 66, 16, 84, 74]. Real-Time [728, 875, 395, 693, 371, 316, 618, 744, 814, 337, 583, 617, 745, 660, 720, 258, 497, 207, 486, 455, 291, 67, 80, 149, 7, 99, 66, 16, 84, 74]. reality [146]. realization [76]. really [38]. reasonable [22]. records [79]. recovery [209]. Reduction [465]. Refactoring [545]. Regional [398]. Related [816, 772, 108]. relationship [87]. Release [679, 692, 755, 783, 873]. Reliability [234, 138]. Reliable [695, 547, 433, 865, 770, 68, 72]. Remote [536]. Removing [115, 557]. rendezvous [66]. Replacing [79]. Report [129, 183, 204, 206, 618, 744, 440, 535, 745, 512, 733, 78, 139, 128, 130, 92]. requeue [475]. Requirements [596, 849, 67]. Research [36, 862]. resource [730]. response [32]. Restrictions [476]. retrospective [210]. Reusable [80, 859, 63, 7, 144, 5, 45, 90]. Reuse [388, 226, 416]. Review [164, 258, 49, 569]. Revision [594, 131]. rewriting [117]. Risk [863]. robust [90]. role [44]. Round [817]. RTFM [788]. RTFM-lang [788]. RTOS [308]. Rules [526]. Runtime [315].

Safe [240, 865]. Safety [354, 596, 816, 814, 571, 772, 849, 859, 291, 56, 108]. Safety- [354]. Safety-Critical [596]. Safety-Related [816]. SAMATE [464]. Scheme [827]. science [18]. Scope [344]. Second [49]. Secure [472, 85, 84]. Security [354]. Security-Critical [354]. see [39]. Seeking [814, 628]. Semantic [198, 208]. Semantics [788, 290]. Seminar [168, 188]. Sensors [47]. Sequence [770]. server [70, 363]. Servers [487]. Session [509, 496, 615, 731, 758, 876, 485, 757, 877, 499, 512, 729, 760, 473, 616, 759, 730, 878, 511, 732]. set [14]. sharing [628]. shelf [108]. should [185]. Shrinks [545]. SIERRADA [2]. Signaling [791]. Simulating [864, 879]. simulation [86]. simulator [107, 696, 132]. Situational [849]. Situations [225]. Sketch [846]. SMP [500]. soccer [696]. Software [138, 155, 100, 638, 697, 709, 682, 708, 433, 847, 440, 836, 514, 512, 799, 548, 670, 770, 513, 345, 248, 556, 569, 74, 773, 73, 56, 29, 185, 695, 106, 176, 52, 472, 144, 5, 120, 45, 68, 72, 108, 116, 16, 408, 211, 148, 606, 619, 36]. software-download [472]. solution [85]. Solve [345]. Some [289, 40, 92]. SonarQube [836]. Source [836, 865]. Sources [151, 163, 177, 190, 199, 214, 237, 135]. Spacecraft [773]. SPARK [837, 274, 584, 619, 774, 802, 682, 683, 761, 838, 770]. SPATS [145]. Specification [3, 629, 45, 198, 208]. specifications [117, 15, 35, 90, 76, 95]. Splitting [70]. sponsoring [408]. Sponsors [200, 215, 238, 250, 260, 268, 276, 300, 309, 282, 293, 317, 324, 331, 346, 356, 365, 373, 380, 389, 399, 409, 417, 426, 434, 442, 449, 457, 609, 622, 632, 641, 663, 686, 700, 736, 748, 764, 819, 777, 793, 804, 829, 840, 851, 867, 881, 187, 652, 673]. Springer [783]. Stack [708]. STAMP [150]. Standard [679, 717, 873, 56, 21, 162, 16, 131, 783]. Standardization [362, 131]. Standardizing [352]. Standards [509, 859, 719, 12]. Static [464, 484, 788, 345]. Status [536, 59]. stealing [628]. step [6]. stimulus [32]. stimulus-response-nets [32]. story [407]. strategies [628]. Strategy [173, 423]. Stream [501]. Strengths [670]. Structure [387, 680]. Structured [813, 95]. Structures [648]. Student [593, 605, 627, 516]. Study [816, 791, 862, 46, 66, 110, 63]. Submitted [679]. Subsetting [559]. Suggestions [501]. Suitability [308]. Summary [615, 731, 758, 876, 757, 877, 729, 760, 616, 759, 730, 878, 732, 100]. Support [787, 837, 708, 498, 720, 13, 11, 12, 109, 150]. supporting [14]. Surprising [31]. Survey [256, 94]. Survivability [849]. Swiss [147]. Switched [790]. Symposium [857]. Syntax [290, 50]. System [811, 860, 524, 465, 638, 814, 849, 256, 708, 173, 863, 498, 799, 398, 743, 21, 52, 46, 113, 107, 10, 407, 132, 1]. Systems [787, 785, 615, 791, 525, 547, 786, 308, 772, 246, 789, 299, 484, 788, 865, 510, 290, 757, 745, 862, 258, 476, 501, 500, 117, 85, 106, 266, 149, 7, 108, 86, 74, 43].

Table [817]. Taking [22, 27, 31, 39, 50]. Task [111, 801, 827]. Task-Based [827]. Tasking [395, 583, 239, 475, 64, 28, 61, 693]. Tasks [670, 500, 66, 79, 161]. Technical [12, 105]. Techniques [226]. Technologies [354]. Technology [36, 423, 694]. Telecare [863]. Template [456]. Ten [172]. test [75, 120, 150]. testability [89]. Thirty [534]. Thread [498]. threaded [105]. Time [728, 875, 395, 693, 465, 371, 316, 487, 618, 744, 814, 547, 337, 583, 659, 789, 847, 617, 745, 498, 660, 720, 258, 497, 207, 486, 455, 80, 111, 82, 149, 7, 99, 113, 66, 16, 84, 291, 694, 74, 67]. Time-Triggered [789]. Timing [786, 708, 150]. TLDacs [109]. Tolerance [548, 159]. tolerant [209]. Tool [772, 708, 111, 120, 11, 12, 150, 58]. Toolkit [197]. Tools [464]. Toolset [787, 145]. Tour [315]. towers [407]. Traffic [799]. Train [398]. transfer [65]. transformation [35]. Transition [292, 78]. translation [846]. transport [212]. trial [21]. trial-use [21]. Triggered [789]. true [31]. Try [706, 692]. Try-and-Beat-Me [706]. Tunnel [100]. type [34, 81]. Types [156, 386, 801, 77, 90].

UK [20, 103, 142, 191, 250, 260, 268, 276, 282, 293, 424, 544, 733, 140, 133, 200, 215, 238, 221, 23, 94, 10]. UML [456, 299, 784]. UNCLE [13]. Understandable [290]. UNIBASE [52]. uniformity [69]. Units [771]. universities [94]. UNIX [105]. Update [130, 874, 362, 838, 48]. updating [19]. Upon [123]. Usage [659, 708, 82]. Use [813, 257, 516, 816, 683, 784, 681, 745, 21, 266, 5, 94, 14, 6, 148]. User [166, 180, 193, 202, 217, 236, 241, 251, 261, 269, 277, 283, 294, 301, 310, 318, 325, 332, 347, 357, 366, 374, 381, 390, 400, 410, 418, 427, 435, 443, 450, 458, 467, 479, 491, 504, 519, 529, 539, 551, 564, 575, 588, 600, 610, 623, 633, 642, 653, 664, 674, 687, 701, 712, 723, 737, 749, 765, 778, 794, 805, 820, 830, 841, 852, 868, 534, 650, 710, 573, 649, 24, 13]. Using [743, 682, 108, 440, 536, 290, 161, 345, 226, 649, 106, 32, 92, 144, 45, 89, 66, 836, 628, 132, 209]. Utilities [486].

V [50]. vacant [225]. Validated [54, 37]. validation [8, 58]. VDM [14]. Vectors [257]. Vendor [125]. Verifiable [143]. Verification [547, 708, 45]. Version [352]. VI [39]. Via [274, 584]. Victorian [825]. viewpoint [120]. virtual [107, 146]. Visibility [387, 680, 39]. Visionary [825]. Vocabularies [813]. Vulnerabilities [560, 606, 619, 561, 558, 535, 557, 556].

warts [115]. water [107]. Way

[593, 605, 627, 274, 584, 692, 696]. Weaving [848]. Web [178, 184, 363]. Wellings [258]. WG9 [594]. Who [92]. Wide [184]. Winners [692]. Wirth [41]. Word [580]. work [266, 628]. work-seeking [628]. work-sharing [628]. work-stealing [628]. Workshop [206, 728, 811, 785, 860, 875, 618, 660, 556, 207, 455, 744, 337, 617]. workstation [53]. World [184]. Worst [465, 847]. Worst-Case [465, 847]. Writing [227]. wrong [61].

XRete [60].

Years [172, 534, 248].

References

Brosgol:1987:DAB

[1] B. M. Brosgol and G. R. Cuthbert. The development of the Ada binding of the Graphical Kernel System. Ada User, 8 (??):3–9, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Arnaudo:1987:SAP

[2] G. Arnaudo and F. Forest. SIERRADA: an Ada package for queueing network modelling. Ada User, 8(??):11–14, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Pitty:1987:LHS

[3] E. B. Pitty, C. Daniels, and C. E. Adams. Languages for hardware specification and design. Ada User, 8(??): 15–20, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Forest:1987:PA

[4] F. Forest and G. Arnaudo. Parallelism and Ada. Ada User, 8(??):21–26, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Gautier:1987:GUA

[5] R. J. Gautier. Guidelines for the use of Ada in reusable software components. Ada User, 8(??):27–32, ???? 1987. CO-DEN ADUSEB. ISSN 0268-652X.

Keenan:1987:RDF

[6] P. Keenan. The re-use of designs as a first step towards the introduction of Ada component re-use. Ada User, 8(??): 33–40, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Elliott:1987:MAA

[7] A. Elliott and R. D. Alexander. Mascot 3 and Ada: an approach to real-time systems based on reusable components. Ada User, 8(??):41–46, ???? 1987. CO-DEN ADUSEB. ISSN 0268-652X.

Chard:1987:AV

[8] A. Chard. Ada validation. Ada User, 8(??):47–51, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Phillips:1987:AE

[9] K. Phillips. Ada evaluation. Ada User, 8(??):53–60, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Marshall:1987:IUM

[10] I. Marshall and R. H. Pierce. Implementation of the UK MOD Ada evaluation system. Ada User, 8(??):61–64, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Lyons:1987:RDT

[11] T. Lyons and M. Tedd. Recent developments in tool support interfaces CAIS and PCTE. Ada User, 8(??):65–72, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Lyons:1987:TOP

[12] T. Lyons and M. Tedd. Technical overview of PCTE and CAIS (tool support interface standards). Ada User, 8(??):73–78, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Kugler:1987:UUI

[13] H.-J. Kugler and C. Unger. UNCLE — user interfaces for Ada support environments. Ada User, 8(??):79–83, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Heerfordt:1987:SAP

[14] H. M. Heerfordt and P. Villadsen. A set of Ada packages supporting the use of VDM for Ada program development. Ada User, 8(??):85–90, ???? 1987. CO-DEN ADUSEB. ISSN 0268-652X.

Cameron:1987:MJN

[15] J. R. Cameron. Mapping JSD network specifications into Ada. Ada User, 8(??): 91–99, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Mearns:1987:SFD

[16] I. Mearns. A standard framework for developing real-time Ada software. Ada User, 8(??):101–106, ???? 1987. CO-DEN ADUSEB. ISSN 0268-652X.

Richmond:1987:OMD

[17] I. Richmond. An overview of managing design (education and data abstraction). Ada User, 8(??):107–110, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Smith:1987:AIC

[18] P. Smith and P. R. Harris. Ada — the implications for the computer science curriculum. Ada User, 8(??):111– 114, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Burhouse:1987:LCA

[19] R. Burhouse and M. Curtis. A local collaborative Ada professional updating programme. Ada User, 8(??):115–120, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Anonymous:1987:AUI

[20] Anonymous. 6th Ada UK International Conference: Ada — the infrastructure. Ada User, 8(1):??, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Boldyreff:1987:ITS

[21] C. Boldyreff. IEEE trial-use standard portable operating system for computer environments — POSIX. Ada User, 8(1):13–14, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Lester:1987:TIAa

[22] K. Lester. Taking issue with Ada. II. Murphy’s law, pre-emption, and being (un)reasonable. Ada User, 8(1):15–16, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Brassington:1987:AAU

[23] R. A. Brassington. Ada activities in UK defence. Ada User, 8(1):31–36, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Kim:1987:AUU

[24] N. Ys. Kim. Ada in the US: a user’s perspective. Ada User, 8(1):37–43, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Syms:1987:AN

[25] T. Syms. Ada in NATO. Ada User, 8(1): 44–47, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Anonymous:1987:CAa

[26] Anonymous. Compilers (Ada). Ada User, 8(2):8–11, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Lester:1987:TIAb

[27] K. Lester. Taking issue with Ada — the cosmology of Ada programs. Ada User, 8(2):13–14, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Burns:1987:ATI

[28] A. Burns, A. M. Lister, and A. J. Wellings. Ada tasking implementation issues. Ada User, 8(2):30–39, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Bishop:1987:APC

[29] J. M. Bishop. Ada profile charts in software development. Ada User, 8(2):40– 45, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Anonymous:1987:CAb

[30] Anonymous. Compilers (Ada). Ada User, 8(3):10–11, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Lester:1987:TIAc

[31] K. Lester. Taking issue with Ada. IV. surprising but true. Ada User, 8(3):15– 16, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Dausmann:1987:RPU

[32] M. Dausmann. Rapid prototyping using stimulus-response-nets and Ada. Ada User, 8(3):33–41, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Pyle:1987:EA

[33] I. C. Pyle. Editors for Ada. Ada User, 8(3):42–45, ???? 1987. CODEN ADUSEB. ISSN 0268-652X.

Hunt:1988:PII

[34] J. Hunt. Packages IMAGES and the IM-AGE data type. Ada User, 9(1):6–13, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Lawton:1988:TJS

[35] J. R. Lawton and N. France. The transformation of JSD specifications into Ada. Ada User, 9(1):29–39, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

McDermid:1988:STR

[36] J. A. McDermid. Software Technology Research Group. Ada User, 9(1):41–46, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Anonymous:1988:CLV

[37] Anonymous. Compilers (list of validated compilers). Ada User, 9(2):57–66, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Marshall:1988:ASR

[38] P. E. Marshall. Ada: is she really a large and difficult lady? Ada User, 9(2):67– 70, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Lester:1988:TIAa

[39] K. Lester. Taking issue with Ada. VI. now you see it, now you don’t (visibility issues). Ada User, 9(2):71–72, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Kermarrec:1988:SEA

[40] Y. Kermarrec. Some experiments with Ada. Ada User, 9(2):79–82, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Forrester:1988:WMC

[41] J. M. Forrester. Wirth meets the countess (Modula-2 to Ada conversion). Ada User, 9(2):83–86, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Anonymous:1988:CIA

[42] Anonymous. Cost implications of Ada. Ada User, 9(2):87–93, ???? 1988. CO-DEN ADUSEB. ISSN 0268-652X.

Wichmann:1988:EAS

[43] B. A. Wichmann. Evaluating Ada systems. Ada User, 9(3):101–102, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Davies:1988:RE

[44] M. Davies. The role of the evaluator. Ada User, 9(3):102–104, ???? 1988. CO-DEN ADUSEB. ISSN 0268-652X.

Hill:1988:FSV

[45] A. Hill. The formal specification and verification of reusable software components using Ada with Asphodel. Ada User, 9(3):113–123, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Foulkes:1988:DCC

[46] R. Foulkes. Design of a command and control system: a MASCOT 3 design study. Ada User, 9(3):124–126, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Gordon:1988:IAA

[47] A. M. Gordon. Introducing Ada at GEC Sensors Ltd. Ada User, 9(3):127–128, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Morron:1988:PU

[48] M. Morron. PCTE update. Ada User, 9(3):129–130, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Dowling:1988:SPI

[49] E. J. Dowling. Second PCTE+ international review. Ada User, 9(3):131–135, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Lester:1988:TIAb

[50] K. Lester. Taking issue with Ada V — Ada syntax, but not Ada. Ada User, 9(4):151–154, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Snowdon:1988:BOI

[51] R. A. Snowdon. A brief overview of the IPSE 2.5 project. Ada User, 9(4):156– 161, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Dittrich:1988:DDS

[52] K. R. Dittrich, W. Gotthard, and P. C. Lockemann. DAMOKLES — the database system for the UNIBASE software engineering environment. Ada User, 9(4):162–172, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Dipper:1988:GWA

[53] H. Dipper and R. Klebe. A graphic workstation for Ada. Ada User, 9(4): 173–178, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Anonymous:1988:VCA

[54] Anonymous. Validated compilers (for Ada, buyer’s guide). Ada User, 9(4): 184–197, ???? 1988. CODEN ADUSEB. ISSN 0268-652X.

Webb:1990:MA

[55] J. T. Webb. MALPAS and Ada. Ada User, 11(1):14–16, ???? 1990. CODEN ADUSEB. ISSN 0268-652X.

Bennett:1990:NIS

[56] P. A. Bennett. Note on IEC software safety standard. Ada User, 11(1):17–20, ???? 1990. CODEN ADUSEB. ISSN 0268-652X.

Wichmann:1990:IA

[57] B. A. Wichmann and S. J. Dawes. Insecurities in Ada. Ada User, 11(1):21–25, ???? 1990. CODEN ADUSEB. ISSN 0268-652X.

Wichmann:1990:LAV

[58] B. A. Wichmann. Low-Ada: an Ada validation tool. Ada User, 11(1):27–32, ???? 1990. CODEN ADUSEB. ISSN 0268-652X.

Anderson:1990:APS

[59] C. Anderson. The Ada 9X project — current status. Ada User, 11(2):67–??, ???? 1990. CODEN ADUSEB. ISSN 0268-652X.

Ardouin:1990:XMI

[60] C. Ardouin, P. Boulle, and M. Menes. XRete: a means to integrate. Ada User, 11(2):88–97, ???? 1990. CODEN ADUSEB. ISSN 0268-652X.

Elsom:1990:WWT

[61] K. C. Elsom. So what’s wrong with tasking? (ada). Ada User, 11(3):128–132, ???? 1990. CODEN ADUSEB. ISSN 0268-652X.

Corbin:1990:FOP

[62] M. J. Corbin and G. F. Butler. A framework for object-oriented programming in Ada. Ada User, 11(3):133–144, ???? 1990. CODEN ADUSEB. ISSN 0268-652X.

Toole:1990:RAC

[63] M. Toole. Reusable Ada components — a case study. Ada User, 11(4):161–164, ???? 1990. CODEN ADUSEB. ISSN 0268-652X.

Baker:1990:OAT

[64] T. Baker. Opening up Ada tasking. Ada User, 11(4):165–169, ???? 1990. CO-DEN ADUSEB. ISSN 0268-652X.

Burns:1990:ATC

[65] A. Burns, A. J. Wellings, and G. L. Davies. Asynchronous transfer of control in Ada 9X. Ada User, 11(4):170–179, ???? 1990. CODEN ADUSEB. ISSN 0268-652X.

Locke:1990:PRA

[66] C. D. Locke, D. R. Vogel, and T. J. Mesler. Predictable real-time avionics design using Ada tasks and rendezvous: a case study. Ada User, 11(4):180–188, ???? 1990. CODEN ADUSEB. ISSN 0268-652X.

Wellings:1990:RRA

[67] A. Wellings. Real-time requirements (for Ada 9X). Ada User, 11(4):189–203, ???? 1990. CODEN ADUSEB. ISSN 0268-652X.

Hill:1991:CPLa

[68] A. D. Hill. The choice of programming language for highly reliable software — a comparison of C and Ada. 1. Ada User, 12(1):11–31, March 1991. CODEN ADUSEB. ISSN 0268-652X.

Kenward:1991:AUI

[69] P. D. Kenward and B. A. Wichmann. Approved uniformity issues (ada). Ada User, 12(1):32–36, March 1991. CODEN ADUSEB. ISSN 0268-652X.

Crowe:1991:SIA

[70] M. K. C. Crowe and P. E. Oldfield. Splitting an incremental Ada environment into server and client. Ada User, 12(1):

37–41, March 1991. CODEN ADUSEB. ISSN 0268-652X.

Riche:1991:FHO

[71] G. J. Riche. Families of high order languages — a rationale. Ada User, 12(1): 42–44, March 1991. CODEN ADUSEB. ISSN 0268-652X.

Hill:1991:CPLb

[72] A. D. Hill. The choice of programming language for highly reliable software — a comparison of C and Ada. 2. Ada User, 12(2):92–103, June 1991. CODEN ADUSEB. ISSN 0268-652X.

Bartholomew-Biggs:1991:NAO

[73] M. C. Bartholomew-Biggs. New algebras for old-numerical software in Ada. Ada User, 12(3):131–136, September 1991. CODEN ADUSEB. ISSN 0268-652X.

Wearing:1991:SDM

[74] A. Wearing. Software development metrics for real-time embedded systems. Ada User, 12(3):137–142, September 1991. CODEN ADUSEB. ISSN 0268-652X.

Gilchrist:1992:GPA

[75] I. Gilchrist and J. Liddiard. A general purpose Ada test harness. Ada User, 13(1):21–26, March 1992. CODEN ADUSEB. ISSN 0268-652X.

Manas:1992:ARL

[76] J. A. Manas, J. Salvachua, M. Veiga, and A. Alvarez. Ada realization of LOTOS specifications. Ada User, 13(1):27–33, March 1992. CODEN ADUSEB. ISSN 0268-652X.

Gauthier:1992:PTC

[77] M. Gauthier. Predefined types considered harmful. Ada User, 13(1):33–43, March 1992. CODEN ADUSEB. ISSN 0268-652X.

Anderson:1992:APR

[78] C. Anderson. Ada 9X project report: transition plan. Ada User, 13(2):51–57, June 1992. CODEN ADUSEB. ISSN 0268-652X.

Locke:1992:RPT

[79] C. D. Locke, T. J. Mesler, and D. R. Vogel. Replacing passive tasks with Ada9X protected records. Ada User, 13(2):58– 62, June 1992. CODEN ADUSEB. ISSN 0268-652X.

Alonso:1992:RRE

[80] A. Alonso and J. A. De La Puente. Reusable real-time executive in Ada: design issues. Ada User, 13(2):63–67, June 1992. CODEN ADUSEB. ISSN 0268-652X.

Losavio:1992:DTB

[81] F. Losavio, A. Matteo, and F. Schlienger. Data type based construction of Ada programs. Ada User, 13(2):68–72, June 1992. CODEN ADUSEB. ISSN 0268-652X.

Burns:1992:MME

[82] A. Burns and J. Wellings. Measuring, monitoring and enforcing CPU execution time usage. Ada User, 13(2):73–78, June 1992. CODEN ADUSEB. ISSN 0268-652X.

Dobbing:1992:EPM

[83] B. J. Dobbing. Experiences with the partitions model. Ada User, 13(2):79–84, June 1992. CODEN ADUSEB. ISSN 0268-652X.

Munch:1992:DSR

[84] P. H. Munch. Development of secure real-time applications in Ada. Ada User, 13(3-4):114–116, September– October 1992. CODEN ADUSEB. ISSN 0268-652X.

Bodilsen:1992:DSS

[85] S. Bodilsen. DDC-I’s solution to secure Ada systems. Ada User, 13(3-4):117– 118, September–October 1992. CODEN ADUSEB. ISSN 0268-652X.

Sherif:1992:ASL

[86] Y. S. Sherif and R. G. Dear. Ada as a simulation language for the analysis of queueing systems. Ada User, 13(3-4): 119–128, September–October 1992. CO-DEN ADUSEB. ISSN 0268-652X.

Longhurst:1993:DAR

[87] D. Longhurst. Defence and Ada — continuing the relationship. Ada User, 14(1):7–10, March 1993. CODEN ADUSEB. ISSN 0268-652X.

Firesmith:1993:POM

[88] D. G. Firesmith. Planning for objectoriented military projects. Ada User, 14(1):11–25, March 1993. CODEN ADUSEB. ISSN 0268-652X.

Liddiard:1993:ATW

[89] J. Liddiard. Achieving testability when using Ada packaging and data hiding methods. Ada User, 14(1):27–32, March 1993. CODEN ADUSEB. ISSN 0268-652X.

Losavio:1993:ARR

[90] F. Losavio. Ada robust and reusable components from algebraic specifications of abstract data types. Ada User, 14(2):55–69, June 1993. CODEN ADUSEB. ISSN 0268-652X.

Moore:1993:IAI

[91] J. W. Moore. The impact of Ada 9X incompatibilities on projects converting from Ada 83. Ada User, 14(2):77–85, June 1993. CODEN ADUSEB. ISSN 0268-652X.

Feldman:1993:WUA

[92] M. B. Feldman. The “who’s using Ada” report — some non-defense Ada applications. Ada User, 14(3):111–115, September 1993. CODEN ADUSEB. ISSN 0268-652X.

Anonymous:1993:AB

[93] Anonymous. Ada bibliography. Ada User, 14(3):117–119, September 1993. CODEN ADUSEB. ISSN 0268-652X.

Glynn:1993:SUA

[94] G. Glynn. Survey of the use of Ada in the universities and colleges in the UK. Ada User, 14(3):121–123, September 1993. CODEN ADUSEB. ISSN 0268-652X.

Priestley:1993:ISA

[95] M. Priestley. Implementing structured algebraic specifications in Ada. Ada User, 14(3):125–134, September 1993. CODEN ADUSEB. ISSN 0268-652X.

Barnes:1993:CBA

[96] John Barnes. Compatibility between Ada 83 and Ada 9X. Ada User, 14 (4):158–160, December 1993. CODEN ADUSEB. ISSN 0268-652X. URL http:

//www.adauk.org.uk/pubs/jbcompat.
htm.

Barnes:1993:OPA

[97] John Barnes. Object-oriented programming in Ada 9X. Ada User, 14(4): 161–163, December 1993. CODEN ADUSEB. ISSN 0268-652X. URL http:

//www.adauk.org.uk/pubs/jboop.htm.

Barnes:1993:POA

[98] John Barnes. Protected objects in Ada 9X. Ada User, 14(4):164–166, December 1993. CODEN ADUSEB. ISSN

0268-652X. URL http://www.adauk.
org.uk/pubs/jbpo.htm.

Firesmith:1993:OAD

[99] D. G. Firesmith. Object-oriented analysis and design for real-time projects. Ada User, 14(4):167–186, December 1993. CODEN ADUSEB. ISSN 0268-652X.

Bennett:1994:SDC

[100] P. A. Bennett. Software development for the Channel Tunnel: a summary. Ada User, 15(1):14–33, March 1994. CODEN ADUSEB. ISSN 0268-652X.

Smart:1994:PPA

[101] John Smart. Packaging policies for automatic code generation. Ada User, 15(1): 34–41, March 1994. CODEN ADUSEB. ISSN 0268-652X.

Barnes:1994:AHL

[102] John Barnes. The Ada 9X hierarchical library. Ada User, 15(1):42–44, March 1994. CODEN ADUSEB. ISSN 0268-652X.

Anonymous:1994:AUI

[103] Anonymous. Ada UK International Conference. Ada User, 15(2):??, June 1994. CODEN ADUSEB. ISSN 0268-652X.

Tokar:1994:EAM

[104] Joyce L. Tokar. The effect of Ada 9X on mature Ada 83 compilers. Ada User, 15(2):55–65, June 1994. CODEN ADUSEB. ISSN 0268-652X.

Larkham:1994:ATL

[105] Adrian Larkham. Ada — the technical language of choice for UNIX threaded applications. Ada User, 15(2):66–76, June 1994. CODEN ADUSEB. ISSN 0268-652X.

Burns:1994:DSS

[106] A. Burns, A. J. Wellings, A. D. Hutcheon, and R. H. Pierce. Dependable software systems using concurrency. Ada User, 15(2):77–85, June 1994. CODEN ADUSEB. ISSN 0268-652X.

Hosseinzaman:1994:AVN

[107] A. Hosseinzaman and A. Bargiela. Ada’s virtual node based water system simulator. Ada User, 15(2):86–94, June 1994. CODEN ADUSEB. ISSN 0268-652X.

Liddiard:1994:UCO

[108] John Liddiard. Using commercial offthe-shelf software in high integrity and safety related systems. Ada User, 15(2): 95–104, June 1994. CODEN ADUSEB. ISSN 0268-652X.

Oliphint:1994:TSA

[109] Clark Oliphint. TLDacs support for ATAC. Ada User, 15(2):105–107, June 1994. CODEN ADUSEB. ISSN 0268-652X.

Smart:1994:IBA

[110] John Smart. Interworking between Ada and C++: a case study. Ada User, 15(2):108–132, June 1994. CODEN ADUSEB. ISSN 0268-652X.

Briggs:1994:TTL

[111] J. S. Briggs, S. D. Jamieson, G. W. Randall, and I. C. Wand. Task time lines as a debugging tool. Ada User, 15(2):133– 150 (or 133–140??), June 1994. CODEN ADUSEB. ISSN 0268-652X.

Grazebrook:1994:RP

[112] Alvery Grazebrook. Rapid prototyping.

Ada User, 15(2):151–152, June 1994. CODEN ADUSEB. ISSN 0268-652X.

Higgon:1994:DAR

[113] Stephen Higgon and B. Hulmgrew. Distributed Ada real time system and debug environment. Ada User, 15(2):153– 159, June 1994. CODEN ADUSEB. ISSN 0268-652X.

Kruchten:1994:MOD

[114] Philippe Kruchten and Mike Tudball.

Mapping an object-oriented design into Ada. Ada User, 15(2):160–176, June 1994. CODEN ADUSEB. ISSN 0268-652X.

Barnes:1994:RAW

[115] John Barnes. Removing Ada’s warts.

Ada User, 15(2):177–182, June 1994. CODEN ADUSEB. ISSN 0268-652X.

Longhurst:1994:CCS

[116] D. W. Longhurst. Counting the cost of software. Ada User, 15(3):189–201, September 1994. CODEN ADUSEB. ISSN 0268-652X.

Acosta:1994:ASA

[117] E. Acosta and F. Losavio. Algebraic specifications into Ada: An approach based on rewriting systems. Ada User, 15(3):203–220, September 1994. CO-DEN ADUSEB. ISSN 0268-652X.

English:1994:FB

[118] J. English. FTP for beginners. Ada User, 15(4):247–249, December 1994. CODEN ADUSEB. ISSN 0268-652X.

Bundgaard:1994:DAC

[119] J. Bundgaard. The design of an Ada 9X Compilation Environment. Ada User, 15 (4):251–262, December 1994. CODEN ADUSEB. ISSN 0268-652X.

Hennell:1994:DVS

[120] M. A. Hennell and J. A. Prudom. A dual viewpoint software test tool. Ada User, 15(4):265–271, December 1994. CODEN ADUSEB. ISSN 0268-652X.

Stuckey:1995:FCI

[121] Richard Stuckey. A fully conformant implementation of ECMA-162. Ada User, 16(2):83–94, June 1995. CODEN AU-JOET. ISSN 0268-652X.

Barnes:19xx:AHL

[122] John Barnes. The Ada 95 hierarchical library. Ada User Journal, ??(??):??, ???? 19xx. CODEN AUJOET. ISSN

                                        1381-6551. URL http://www.adauk.
                                        org.uk/pubs/jblib.htm.

Barnes:19xx:YCD

[123] John Barnes. You can depend upon Ada. Ada User Journal, ??(??):??, ???? 19xx. CODEN AUJOET. ISSN 1381-

                                        6551. URL http://www.adauk.org.
                                        uk/pubs/jbdepend.htm.

Anonymous:1995:Ea

[124] Anonymous. Editorial. Ada User Journal, 16(1):3–??, March 1995. CODEN AUJOET. ISSN 0268-652X.

Anonymous:1995:VN

[125] Anonymous. Vendor notes. Ada User Journal, 16(1):5–??, March 1995. CO-DEN AUJOET. ISSN 0268-652X.

Anonymous:1995:Na

[126] Anonymous. News. Ada User Journal, 16(1):7–??, March 1995. CODEN AU-JOET. ISSN 0268-652X.

Anonymous:1995:FEa

[127] Anonymous. Forthcoming events. Ada User Journal, 16(1):16–??, March 1995. CODEN AUJOET. ISSN 0268-652X.

Anonymous:1995:CR

[128] Anonymous. Conference report. Ada User Journal, 16(1):22–??, March 1995. CODEN AUJOET. ISSN 0268-652X.

Anonymous:1995:RIA

[129] Anonymous. Report International Ada Liaison Committee. Ada User Journal, 16(1):25–??, March 1995. CODEN AU-JOET. ISSN 0268-652X.

Anonymous:1995:UGR

[130] Anonymous. Update GNAT report. Ada User Journal, 16(1):27–??, March 1995. CODEN AUJOET. ISSN 0268-652X.

Anonymous:1995:SAL

[131] Anonymous. Standard Ada language revision achieves international standardization. Ada User Journal, 16(1):29–??, March 1995. CODEN AUJOET. ISSN 0268-652X.

Woodford:1995:CQS

[132] S. Woodford. Creating a queuing system simulator using Ada. Ada User Journal, 16(1):32–39, March 1995. CODEN AU-JOET. ISSN 0268-652X.

Anonymous:1995:AUM

[133] Anonymous. Ada UK marketing committee. Ada User Journal, 16(1):40–??, March 1995. CODEN AUJOET. ISSN 0268-652X.

Anonymous:1995:FCA

[134] Anonymous. Final changes to Ada 95. Ada User Journal, 16(1):42–??, March 1995. CODEN AUJOET. ISSN 0268-652X.

Anonymous:1995:AIS

[135] Anonymous. Ada information sources. Ada User Journal, 16(1):47–??, March 1995. CODEN AUJOET. ISSN 0268-652X.

Anonymous:1995:Eb

[136] Anonymous. Editorial. Ada User Journal, 16(2):50–??, June 1995. CODEN AUJOET. ISSN 0268-652X.

Anonymous:1995:Nb

[137] Anonymous. News. Ada User Journal, 16(2):53–??, June 1995. CODEN AU-JOET. ISSN 0268-652X.

Anonymous:1995:SRM

[138] Anonymous. Software reliability and metrics club. Ada User Journal, 16(2): 61–??, June 1995. CODEN AUJOET. ISSN 0268-652X.

Anonymous:1995:AR

[139] Anonymous. Ada 9X report. Ada User Journal, 16(2):63–??, June 1995. CO-DEN AUJOET. ISSN 0268-652X.

Anonymous:1995:AUA

[140] Anonymous. Ada UK academic affiliates. Ada User Journal, 16(2):67–??, June 1995. CODEN AUJOET. ISSN 0268-652X.

Bazex:1995:APT

[141] Pierre Bazex, Thierry Millan, and Frederic Mulatero. Ada persistence through an OODBMS O2. Ada User Journal, 16(2):71–82, June 1995. CODEN AU-JOET. ISSN 0268-652X.

Anonymous:1995:AUC

[142] Anonymous. 1995 ADA UK Conference.

Ada User Journal, 16(3):??, September 1995. CODEN AUJOET. ISSN 0268-652X.

Pyle:1995:VAP

[143] I. C. Pyle and F. W. Long. Verifiable Ada packages. Ada User Journal, 16(3): 103–109, September 1995. CODEN AU-JOET. ISSN 0268-652X.

Gargaro:1995:TRS

[144] Anthony Gargaro and A. Spencer Peterson. Towards reusable software architectures using distributed objects. Ada User Journal, 16(3):110–122, September 1995. CODEN AUJOET. ISSN 0268-652X.

Chapman:1995:SNT

[145] R. Chapman, A. Burns, and A. Wellings.

SPATS — a new toolset for highintegrity Ada development. Ada User Journal, 16(3):123–131, September 1995. CODEN AUJOET. ISSN 0268-652X.

McAllister:1995:IAV

[146] Dave McAllister. The impact of Ada on virtual reality. Ada User Journal, 16(3): 132–135, September 1995. CODEN AU-JOET. ISSN 0268-652X.

Paterson:1995:SPA

[147] A. Paterson. Swiss Postbank: Ada in a commercial application. Ada User Journal, 16(3):136–150, September 1995. CODEN AUJOET. ISSN 0268-652X.

Vicary:1995:ULS

[148] S. J. Vicary. The use of legacy software in an Ada environment. Ada User Journal, 16(3):151–162, September 1995. CODEN AUJOET. ISSN 0268-652X.

Chastek:1995:THC

[149] Gary Chastek and Anthony Gargaro. Towards heterogeneous communication for POSIX real-time distributed systems. Ada User Journal, 16(3):163–176, September 1995. CODEN AUJOET. ISSN 0268-652X.

Pierce:1995:STS

[150] R. H. Pierce. STAMP: a tool to support timing test generation. Ada User Journal, 16(3):177–179, September 1995. CODEN AUJOET. ISSN 0268-652X.

Anonymous:1995:ASI

[151] Anonymous. Ada sources of information. Ada User Journal, 16(4):??, December 1995. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1995:Ec

[152] Anonymous. Editorial. Ada User Journal, 16(4):??, December 1995. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1995:FEb

[153] Anonymous. Forthcoming events. Ada User Journal, 16(4):??, December 1995. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1995:Nc

[154] Anonymous. News. Ada User Journal, 16(4):??, December 1995. CODEN AU-JOET. ISSN 1381-6551.

Bamberger:1995:SPP

[155] Judy Bamberger and James Hook. Software process practicum: Lessons in software quality and leadership. Ada User Journal, 16(4):205–211, December 1995. CODEN AUJOET. ISSN 0268-652X.

Barnes:1995:ATA

[156] John Barnes. Access types in Ada 95.

Ada User Journal, 16(4):212–214, December 1995. CODEN AUJOET. ISSN

0268-652X. URL http://www.adauk.
org.uk/pubs/jbaccess.htm.

Barnes:1995:APE

[157] John Barnes. The Ada 95 predefined environment. Ada User Journal, 16(4): 215–219, December 1995. CODEN AU-JOET. ISSN 0268-652X. URL http:

//www.adauk.org.uk/pubs/jbpredef.
htm.

Barnes:1995:EA

[158] John Barnes. Exceptions in Ada 95. Ada User Journal, 16(4):220–223, December 1995. CODEN AUJOET. ISSN 0268-

652X. URL http://www.adauk.org.
uk/pubs/jbexcep.htm.

Kermarrec:1995:IEF

[159] Yvon Kermarrec, Laurent Nana Tchamnda, and Laurent Pautet. Implementing an efficient fault tolerance mechanism in Ada 9X: an early experiment with GNAT. Ada User Journal, 16(4): 224–228, December 1995. CODEN AU-JOET. ISSN 0268-652X.

Zeigler:1996:CDC

[160] Stephen F. Zeigler. Comparing development costs of C and Ada. Ada User Journal, 17(??):??, ???? 1996. CODEN AUJOET. ISSN 1381-6551. URL http:

                                        //www.adauk.org.uk/pubs/zeigler.
                                        htm.

Sanden:1996:UTC

[161] B. I. Sanden. Using tasks to capture problem concurrency. Ada User Journal, 17(1):25–36, March 1996. CODEN AUJOET. ISSN 1381-6551.

DeVroey:1996:MFE

[162] S. De Vroey. My first experience with Ada 83: implementing the standard backpropagation algorithm. Ada User Journal, 17(1):37–45, March 1996. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:1996:ASIa

[163] Anonymous. Ada sources of information. Ada User Journal, 17(2):??, June 1996. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1996:BR

[164] Anonymous. Book review. Ada User Journal, 17(2):??, June 1996. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1996:E

[165] Anonymous. Editorial. Ada User Journal, 17(2):??, June 1996. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:1996:EPAa

[166] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 17(2): ??, June 1996. CODEN AUJOET. ISSN

                                        1381-6551. URL http://www.adauk.
                                        org.uk/pubs/jouedpol.htm.

Anonymous:1996:FEa

[167] Anonymous. Forthcoming events. Ada User Journal, 17(2):??, June 1996. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:1996:IAS

[168] Anonymous. Invitation to Ada 95 seminar. Ada User Journal, 17(2):??, June 1996. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1996:MYN

[169] Anonymous. A message from your new Chairman. Ada User Journal, 17(2):??, June 1996. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1996:Na

[170] Anonymous. News. Ada User Journal, 17(2):??, June 1996. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1996:NMC

[171] Anonymous. News from the Marketing Committee. Ada User Journal, 17(2):??, June 1996. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1996:TYT

[172] Anonymous. Ten years and ten days.

Ada User Journal, 17(2):??, June 1996. CODEN AUJOET. ISSN 1381-6551.

MenendezdeLlanoRozas:1996:ADS

[173] R. Menendez de Llano Rozas and J. A. G. Monasterio. An alternative distribution strategy of an Ada program in a multicomputer system. Ada User Journal, 17(2):72–88, June 1996. CO-DEN AUJOET. ISSN 1381-6551.

Ellis:1996:OOD

[174] Rod Ellis. Object oriented design patterns in Ada 95. Ada User Journal, 17 (2):89–104, June 1996. CODEN AU-JOET. ISSN 1381-6551.

Smith:1996:OPA

[175] M. Smith. OO-programming in Ada. Ada User Journal, 17(3):125–131, September 1996. CODEN AUJOET. ISSN 1381-6551.

Culwin:1996:BLS

[176] F. Culwin. Beyond the LAW, a software development pedagogy for the 21st century. Ada User Journal, 17(3):132–142, September 1996. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1996:ASIb

[177] Anonymous. Ada sources of information. Ada User Journal, 17(4):??, December 1996. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1996:AW

[178] Anonymous. AdaUK on the Web. Ada User Journal, 17(4):??, December 1996. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1996:EP

[179] Anonymous. Editorial page. Ada User Journal, 17(4):??, December 1996. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:1996:EPAb

[180] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 17(4): ??, December 1996. CODEN AUJOET. ISSN 1381-6551. URL http://www.

                                        adauk.org.uk/pubs/jouedpol.htm.

Anonymous:1996:FEb

[181] Anonymous. Forthcoming events. Ada User Journal, 17(4):??, December 1996. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1996:Nb

[182] Anonymous. News. Ada User Journal, 17(4):??, December 1996. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:1996:RMC

[183] Anonymous. Report from the Marketing Committee. Ada User Journal, 17(4): ??, December 1996. CODEN AUJOET. ISSN 1381-6551.

Briggs:1996:AWW

[184] Jim Briggs. Ada on the World Wide Web. Ada User Journal, 17(4):162–164, December 1996. CODEN AUJOET. ISSN 0268-652X.

Briggs:1996:DSP

[185] Jim Briggs. DoD software policy — should Ada be mandated? Ada User Journal, 17(4):165–167, December 1996. CODEN AUJOET. ISSN 0268-652X.

Barnes:1996:WDK

[186] J. Barnes. We don’t know nothing [programming languages]. Ada User Journal, 17(4):168–171, December 1996. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:1996:MS

[187] Anonymous. Meet the sponsors. Ada User Journal, 17(4):172–176, December 1996. CODEN AUJOET. ISSN 0268-652X.

Anonymous:1997:ABS

[188] Anonymous. Ada briefing seminar. Ada User Journal, 18(1):??, March 1997. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1997:AEB

[189] Anonymous. Ada Europe board. Ada User Journal, 18(1):??, March 1997. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1997:ASIa

[190] Anonymous. Ada sources of information. Ada User Journal, 18(1):??, March 1997. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1997:AUA

[191] Anonymous. The Ada UK Academic Committee. Ada User Journal, 18(1):??, March 1997. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1997:Ea

[192] Anonymous. Editorial. Ada User Journal, 18(1):??, March 1997. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:1997:EPAa

[193] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 18 (1):??, March 1997. CODEN AUJOET. ISSN 1381-6551. URL http://www.

                                        adauk.org.uk/pubs/jouedpol.htm.

Anonymous:1997:FEa

[194] Anonymous. Forthcoming events. Ada User Journal, 18(1):??, March 1997. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1997:LE

[195] Anonymous. Letters to the Editor. Ada User Journal, 18(1):??, March 1997. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1997:Na

[196] Anonymous. News. Ada User Journal, 18(1):??, March 1997. CODEN AU-JOET. ISSN 1381-6551.

Betz:1997:EAC

[197] H. Betz, B. Holzmuller, and R. Koschke. Experiences in adjusting a compiler toolkit to generate Ada 95 code. Ada User Journal, 18(1):29–40, March 1997. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1997:ASIb

[198] Anonymous. Ada Semantic Interface Specification frequently asked questions (ASISFAQ). Ada User Journal, 18(2): ??, June 1997. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1997:ASIc

[199] Anonymous. Ada sources of information. Ada User Journal, 18(2):??, June 1997. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1997:AUS

[200] Anonymous. Ada UK sponsors. Ada User Journal, 18(2):??, June 1997. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:1997:Eb

[201] Anonymous. Editorial. Ada User Journal, 18(2):??, June 1997. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:1997:EPAb

[202] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 18(2): ??, June 1997. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1997:FEb

[203] Anonymous. Forthcoming events. Ada User Journal, 18(2):??, June 1997. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:1997:MRA

[204] Anonymous. Meeting report — ASISWG/ASISRG. Ada User Journal, 18(2):??, June 1997. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1997:Nb

[205] Anonymous. News. Ada User Journal, 18(2):??, June 1997. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1997:WRI

[206] Anonymous. Workshop report — IRTAW8. Ada User Journal, 18(2):??, June 1997. CODEN AUJOET. ISSN 1381-6551.

Wellings:1997:EIR

[207] Andy Wellings and Alan Burns. Eight International Real-Time Ada Workshop (IRTAW8). Ada User Journal, 18(2):66– 72, June 1997. CODEN AUJOET. ISSN 0268-652X.

Anonymous:1997:ASId

[208] Anonymous. Ada Semantic Interface Specification (ASIS) frequently asked questions. Ada User Journal, 18(2):75– 77, June 1997. CODEN AUJOET. ISSN 0268-652X.

Zweiacker:1997:MCA

[209] M. Zweiacker. Making CORBA applications fault tolerant using checkpointing and recovery. Ada User Journal, 18(2): 78–91, June 1997. CODEN AUJOET. ISSN 1381-6551.

Ward:1997:RRE

[210] W. A. Ward, Jr. A retrospective on the Rational Environment. Ada User Journal, 18(3):121–143, September 1997. CODEN AUJOET. ISSN 1381-6551.

Simpson:1997:ASE

[211] D. Simpson. Ada and software engineering education. Ada User Journal, 18(3): 144–147, September 1997. CODEN AU-JOET. ISSN 1381-6551.

GonzalezBarahona:1997:LLF

[212] Jesus M. Gonzalez Barahona, Pedro de las Heras Quiros, and Jose Centeno Gonzalez. Lower Layer: a family of interfaces to transport communication protocols. Ada User Journal, 18(4):171–189, December 1997. CODEN AUJOET. ISSN 0268-652X.

Anonymous:1998:APC

[213] Anonymous. Ada programming competition. Ada User Journal, 19(1):??, April 1998. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1998:ASIa

[214] Anonymous. Ada sources of information. Ada User Journal, 19(1):??, April 1998. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1998:AUSa

[215] Anonymous. Ada UK sponsors. Ada User Journal, 19(1):??, April 1998. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:1998:E

[216] Anonymous. Editorial. Ada User Journal, 19(1):??, April 1998. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:1998:EPAa

[217] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 19(1): ??, April 1998. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1998:FEa

[218] Anonymous. Forthcoming Events. Ada User Journal, 19(1):??, April 1998. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:1998:IAE

[219] Anonymous. Introduction to Ada Europe. Ada User Journal, 19(1):??, April 1998. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1998:KPF

[220] Anonymous. Kissing the polymorphic frog. Ada User Journal, 19(1):??, April 1998. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1998:MCA

[221] Anonymous. A message from the Chairman of Ada UK. Ada User Journal, 19 (1):??, April 1998. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1998:MPA

[222] Anonymous. A message from the President of Ada–Europe. Ada User Journal, 19(1):??, April 1998. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1998:Na

[223] Anonymous. News. Ada User Journal, 19(1):??, April 1998. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1998:P

[224] Anonymous. Papers. Ada User Journal, 19(1):??, April 1998. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1998:SV

[225] Anonymous. Situations vacant. Ada User Journal, 19(1):??, April 1998. CO-DEN AUJOET. ISSN 1381-6551.

Yu:1998:UOT

[226] Huiming Yu. Using object-oriented techniques to develop an Ada 95 Reuse Library. Ada User Journal, 19(1):43–56, April 1998. CODEN AUJOET. ISSN 0268-652X.

Westley:1998:WJA

[227] Terry J. Westley. Writing Java applets in Ada: a beginner’s guide. Ada User Journal, 19(1):57–82, April 1998. CO-DEN AUJOET. ISSN 0268-652X.

Briggs:1998:E

[228] Jim Briggs. Editorial. Ada User Journal, 19(2):89–??, ???? 1998. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1998:Nb

[229] Anonymous. News. Ada User Journal, 19(2):90–??, ???? 1998. CODEN AU-JOET. ISSN 1381-6551.

Barnes:1998:AEG

[230] John Barnes. The Ada–Europe General Assembly. Ada User Journal, 19(2):110– ??, ???? 1998. CODEN AUJOET. ISSN 1381-6551.

Barnes:1998:AEC

[231] John Barnes. The 1998 Ada–Europe Conference. Ada User Journal, 19(2): 114–??, ???? 1998. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1998:CC

[232] Anonymous. Conference calendar. Ada User Journal, 19(2):118–??, ???? 1998. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1998:FEb

[233] Anonymous. Forthcoming events. Ada User Journal, 19(2):121–??, ???? 1998. CODEN AUJOET. ISSN 1381-6551.

Riehle:1998:RDL

[234] Richard Riehle. Reliability — does language matter? Ada User Journal, 19(2):

128–??, ???? 1998. CODEN AUJOET. ISSN 1381-6551.

Brosgol:1998:WAR

[235] Ben Brosgol. Why Ada rather than Java for academia? Ada User Journal, 19(2): 135–??, ???? 1998. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1998:EPAb

[236] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 19(2): 137–??, ???? 1998. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1998:ASIb

[237] Anonymous. Ada sources of information. Ada User Journal, 19(2):138–??, ???? 1998. CODEN AUJOET. ISSN 1381-6551.

Anonymous:1998:AUSb

[238] Anonymous. Ada UK sponsors. Ada User Journal, 19(2):139–??, ???? 1998. CODEN AUJOET. ISSN 1381-6551.

Nonaka:1999:TDD

[239] Yusuke Nonaka, Jingde Cheng, and Kazuo Ushijima. A tasking deadlock detector for Ada 95 programs. Ada User Journal, 20(1):79–92, April 1999. CODEN AUJOET. ISSN 1381-

6551. URL http://citeseerx.ist.
psu.edu/viewdoc/download?doi=10.
1.1.42.247&rep=rep1&type=pdf.

Grein:1999:SP

[240] Christoph Karl Walter Grein. Safe pointers. Ada User Journal, 20(2): ??, July 1999. CODEN AUJOET. ISSN 1381-6551. URL http://christ-

usch-grein.homepage.t-online.de/
Ada/Safe_Pointers.html.

Anonymous:2001:EPAa

[241] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 22(1): 2–??, March 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:Ea

[242] Anonymous. Editorial. Ada User Journal, 22(1):3–??, March 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:Na

[243] Anonymous. News. Ada User Journal, 22(1):5–??, March 2001. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2001:CCa

[244] Anonymous. Conference calendar. Ada User Journal, 22(1):24–??, March 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:FEa

[245] Anonymous. Forthcoming events. Ada User Journal, 22(1):31–??, March 2001. CODEN AUJOET. ISSN 1381-6551.

Humphris:2001:IAD

[246] David Humphris. Integrating Ada into a distributed systems environment. Ada User Journal, 22(1):36–??, March 2001. CODEN AUJOET. ISSN 1381-6551.

Sanchez:2001:ADP

[247] Fernando Sánchez, Roberto Rodríguez, Juan M. Murillo, and Juan Hernández. Adaptability of distribution protocols in object-oriented languages. Ada User Journal, 22(1):44–??, March 2001. CO-DEN AUJOET. ISSN 1381-6551.

Thombs:2001:MIM

[248] David Thombs. The major issues for MOD software procurement over the next five years. Ada User Journal, 22(1): 52–??, March 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:AEAa

[249] Anonymous. Ada–Europe associate members (national Ada organisations). Ada User Journal, 22(1):59–??, March 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:AUSa

[250] Anonymous. Ada UK 2001 sponsors. Ada User Journal, 22(1):60–??, March 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:EPAb

[251] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 22(2): 62–??, June 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:Eb

[252] Anonymous. Editorial. Ada User Journal, 22(2):63–??, June 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:Nb

[253] Anonymous. News. Ada User Journal, 22(2):65–??, June 2001. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2001:CCb

[254] Anonymous. Conference calendar. Ada User Journal, 22(2):104–??, June 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:FEb

[255] Anonymous. Forthcoming events. Ada User Journal, 22(2):110–??, June 2001. CODEN AUJOET. ISSN 1381-6551.

Gilchrist:2001:CAL

[256] Ian Gilchrist. The contribution of the Ada language to system development: a market survey. Ada User Journal, 22(2): 118–??, June 2001. CODEN AUJOET. ISSN 1381-6551.

Barnes:2001:DYU

[257] John Barnes. Do you use generic arrays and vectors? — ISO/IEC 13813:1998. Ada User Journal, 22(2):124–??, June 2001. CODEN AUJOET. ISSN 1381-6551.

Rogers:2001:BRR

[258] Pat Rogers. Book review: Real-Time Systems and Programming Languages by Alan Burns and Andy Wellings. Ada User Journal, 22(2):126–??, June 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:AEAb

[259] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 22(2):127–??, June 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:AUSb

[260] Anonymous. Ada UK 2001 sponsors.

Ada User Journal, 22(2):128–??, June 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:EPAc

[261] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 22(3): 130–??, September 2001. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2001:Ec

[262] Anonymous. Editorial. Ada User Journal, 22(3):131–??, September 2001. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2001:Nc

[263] Anonymous. News. Ada User Journal, 22(3):133–??, September 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:CCc

[264] Anonymous. Conference calendar. Ada User Journal, 22(3):171–??, September 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:FEc

[265] Anonymous. Forthcoming events. Ada User Journal, 22(3):177–??, September 2001. CODEN AUJOET. ISSN 1381-6551.

Burns:2001:GUA

[266] Alan Burns. Guide for the use of the Ada Ravenscar Profile in high integrity systems — the work of the HRG. Ada User Journal, 22(3):182–??, September 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:AEAc

[267] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 22(3):189– ??, September 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:AUSc

[268] Anonymous. Ada UK 2001 sponsors. Ada User Journal, 22(3):190–??, September 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:EPAd

[269] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 22(4): 192–??, December 2001. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2001:Ed

[270] Anonymous. Editorial. Ada User Journal, 22(4):193–??, December 2001. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2001:Nd

[271] Anonymous. News. Ada User Journal, 22(4):195–??, December 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:CCd

[272] Anonymous. Conference calendar. Ada User Journal, 22(4):233–??, December 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:FEd

[273] Anonymous. Forthcoming events. Ada User Journal, 22(4):239–??, December 2001. CODEN AUJOET. ISSN 1381-6551.

Barnes:2001:SWC

[274] John Barnes. The SPARK way to correctness is via abstraction. Ada User Journal, 22(4):244–??, December 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:AEAd

[275] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 22(4):255–??, December 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2001:AUSd

[276] Anonymous. Ada UK 2001 sponsors.

Ada User Journal, 22(4):256–??, December 2001. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:EPAa

[277] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 23(1):

2–??, March 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:Ea

[278] Anonymous. Editorial. Ada User Journal, 23(1):3–??, March 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:Na

[279] Anonymous. News. Ada User Journal, 23(1):5–??, March 2002. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2002:CCa

[280] Anonymous. Conference calendar. Ada User Journal, 23(1):44–??, March 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:FEa

[281] Anonymous. Forthcoming events. Ada User Journal, 23(1):51–??, March 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:AUSa

[282] Anonymous. Ada UK 2002 sponsors. Ada User Journal, 23(1):64–??, March 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:EPAb

[283] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 23(2): 66–??, June 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:Eb

[284] Anonymous. Editorial. Ada User Journal, 23(2):67–??, June 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:Nb

[285] Anonymous. News. Ada User Journal, 23(2):69–??, June 2002. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2002:CCb

[286] Anonymous. Conference calendar. Ada User Journal, 23(2):83–??, June 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:FEb

[287] Anonymous. Forthcoming events. Ada User Journal, 23(2):91–??, June 2002. CODEN AUJOET. ISSN 1381-6551.

Leroy:2002:CA

[288] Pascal Leroy. Call for APIs. Ada User Journal, 23(2):96–??, June 2002. CO-DEN AUJOET. ISSN 1381-6551.

Barnes:2002:SII

[289] John Barnes. Some impressions from IRTAW 11. Ada User Journal, 23(2): 97–??, June 2002. CODEN AUJOET. ISSN 1381-6551.

Ogren:2002:UAS

[290] Ingmar Ögren. Using Ada’s syntax and semantics for understandable systems engineering. Ada User Journal, 23(2): 100–??, June 2002. CODEN AUJOET. ISSN 1381-6551.

Sanden:2002:RTP

[291] Bo I. Sandén. Real-time programming safety in Java and Ada. Ada User Journal, 23(2):105–??, June 2002. CODEN AUJOET. ISSN 1381-6551.

Cousins:2002:TLP

[292] Jeff Cousins. Transition of a large project from Ada 83 to Ada 95. Ada User Journal, 23(2):114–??, June 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:AUSb

[293] Anonymous. Ada UK 2002 sponsors.

Ada User Journal, 23(2):119–??, June 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:EPAc

[294] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 23(3): 122–??, September 2002. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2002:Ec

[295] Anonymous. Editorial. Ada User Journal, 23(3):123–??, September 2002. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2002:Nc

[296] Anonymous. News. Ada User Journal, 23(3):125–??, September 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:CCc

[297] Anonymous. Conference calendar. Ada User Journal, 23(3):160–??, September 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:FEc

[298] Anonymous. Forthcoming events. Ada User Journal, 23(3):167–??, September 2002. CODEN AUJOET. ISSN 1381-6551.

Kersten:2002:CUD

[299] Michael Kersten, Joerg Matthes, Christian Fouda Manga, Stephan Zipser, and Hubert B. Keller. Customizing UML for the development of distributed reactive systems and Ada 95 code generation. Ada User Journal, 23(3):175– ??, September 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:AESa

[300] Anonymous. Ada–Europe 2002 sponsors. Ada User Journal, 23(3):184–??, September 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:EPAd

[301] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 23(4): 186–??, December 2002. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2002:Ed

[302] Anonymous. Editorial. Ada User Journal, 23(4):187–??, December 2002. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2002:Nd

[303] Anonymous. News. Ada User Journal, 23(4):189–??, December 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:CCd

[304] Anonymous. Conference calendar. Ada User Journal, 23(4):219–??, December 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:FEd

[305] Anonymous. Forthcoming events. Ada User Journal, 23(4):225–??, December 2002. CODEN AUJOET. ISSN 1381-6551.

Leroy:2002:IA

[306] Pascal Leroy. An invitation to Ada 2005.

Ada User Journal, 23(4):230–??, December 2002. CODEN AUJOET. ISSN 1381-6551.

Santos:2002:ECA

[307] Luìs Santos. Evolutionary computation in Ada 95: a genetic algorithm approach. Ada User Journal, 23(4):239– ??, December 2002. CODEN AUJOET. ISSN 1381-6551.

Dobbing:2002:CRS

[308] Brian Dobbing and Terry Totten. COTS RTOS suitability for high integrity systems. Ada User Journal, 23(4):245– ??, December 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2002:AESb

[309] Anonymous. Ada–Europe 2002 sponsors. Ada User Journal, 23(4):248–??, December 2002. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2003:EPAa

[310] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 24(1): 2–??, March 2003. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2003:Ea

[311] Anonymous. Editorial. Ada User Journal, 24(1):3–??, March 2003. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2003:Na

[312] Anonymous. News. Ada User Journal, 24(1):4–??, March 2003. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2003:CCa

[313] Anonymous. Conference calendar. Ada User Journal, 24(1):42–??, March 2003. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2003:FEa

[314] Anonymous. Forthcoming events. Ada User Journal, 24(1):50–??, March 2003. CODEN AUJOET. ISSN 1381-6551.

Miranda:2003:GTG

[315] Javier Miranda. A guided tour to the GNAT runtime. Ada User Journal, 24 (1):55–??, March 2003. CODEN AU-JOET. ISSN 1381-6551.

Burns:2003:IRT

[316] Alan Burns. Increasing the real-time expressive power of Ada 95. Ada User Journal, 24(1):60–??, March 2003. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2003:AESa

[317] Anonymous. Ada–Europe 2002 sponsors. Ada User Journal, 24(1):64–??, March 2003. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2003:EPAb

[318] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 24(2): 66–??, June 2003. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2003:Eb

[319] Anonymous. Editorial. Ada User Journal, 24(2):67–??, June 2003. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2003:Nb

[320] Anonymous. News. Ada User Journal, 24(2):68–??, June 2003. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2003:CCb

[321] Anonymous. Conference calendar. Ada User Journal, 24(2):109–??, June 2003. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2003:FEb

[322] Anonymous. Forthcoming events. Ada User Journal, 24(2):114–??, June 2003. CODEN AUJOET. ISSN 1381-6551.

Taft:2003:OOP

[323] S Tucker Taft. Object-oriented programming enhancements in Ada 200Y. Ada User Journal, 24(2):119–??, June 2003. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2003:AESb

[324] Anonymous. Ada–Europe 2002 sponsors. Ada User Journal, 24(2):128–??, June 2003. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2003:EPAc

[325] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 24(3): 130–??, September 2003. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2003:Ec

[326] Anonymous. Editorial. Ada User Journal, 24(3):131–??, September 2003. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2003:Nc

[327] Anonymous. News. Ada User Journal, 24(3):132–??, September 2003. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2003:CCc

[328] Anonymous. Conference calendar. Ada User Journal, 24(3):173–??, September 2003. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2003:FEc

[329] Anonymous. Forthcoming events. Ada User Journal, 24(3):179–??, September 2003. CODEN AUJOET. ISSN 1381-6551.

Barnes:2003:ANP

[330] John Barnes. Additional numeric packages for Ada 2005. Ada User Journal, 24(3):185–??, September 2003. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2003:AESc

[331] Anonymous. Ada–Europe 2003 sponsors. Ada User Journal, 24(3):192–??, September 2003. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2003:EPAd

[332] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 24(1): 194–??, December 2003. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2003:Ed

[333] Anonymous. Editorial. Ada User Journal, 24(1):195–??, December 2003. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2003:Nd

[334] Anonymous. News. Ada User Journal, 24(1):196–??, December 2003. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2003:CCd

[335] Anonymous. Conference calendar. Ada User Journal, 24(1):230–??, December 2003. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2003:FEd

[336] Anonymous. Forthcoming events. Ada User Journal, 24(1):238–??, December 2003. CODEN AUJOET. ISSN 1381-6551.

Dobbing:2003:IRT

[337] Brian Dobbing, John Barnes, and Miguel Pinho. 12th International Real-Time Ada Workshop. Ada User Journal, 24(1):242–??, December 2003. CODEN AUJOET. ISSN 1381-6551.

Masmano:2003:MFA

[338] Miguel Masmano. My first Ada–Europe conference. Ada User Journal, 24(1): 246–??, December 2003. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2004:Na

[339] Anonymous. News. Ada User Journal, 25(1):1–??, March 2004. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2004:AEG

[340] Anonymous. Ada–Europe general assembly 2004 agenda. Ada User Journal, 25(1):23–??, March 2004. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2004:PMA

[341] Anonymous. Proposed motions for the Ada–Europe general assembly. Ada User Journal, 25(1):24–??, March 2004. CO-DEN AUJOET. ISSN 1381-6551.

WG9:2004:IAP

[342] ISO WG9. Instructions to the ARG for preparation of the amendment to ISO/ IEC 8652. Ada User Journal, 25(1): 27–??, March 2004. CODEN AUJOET. ISSN 1381-6551.

Moore:2004:CIA

[343] James W. Moore. Comments on: Instructions to the ARG for preparation of the amendment to ISO/IEC 8652. Ada User Journal, 25(1):28–??, March 2004. CODEN AUJOET. ISSN 1381-6551.

Leroy:2004:PDS

[344] Pascal Leroy. Proposal for defining scope of amendment to ISO/IEC 8652:1995. Ada User Journal, 25(1):29–??, March 2004. CODEN AUJOET. ISSN 1381-6551.

Taft:2004:FSB

[345] S. Tucker Taft. Fixing software before it breaks: Using static analysis to help solve the software quality quagmire. Ada User Journal, 25(1):33–??, March 2004. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2004:AESa

[346] Anonymous. Ada–Europe 2004 sponsors. Ada User Journal, 25(1):38–??, March 2004. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2004:EPAa

[347] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 25(2): 40–??, June 2004. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2004:Ea

[348] Anonymous. Editorial. Ada User Journal, 25(2):41–??, June 2004. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2004:Nb

[349] Anonymous. News. Ada User Journal, 25(2):43–??, June 2004. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2004:CCa

[350] Anonymous. Conference calendar. Ada User Journal, 25(2):77–??, June 2004. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2004:FEa

[351] Anonymous. Forthcoming events. Ada User Journal, 25(2):85–??, June 2004. CODEN AUJOET. ISSN 1381-6551.

Moore:2004:SNV

[352] James W. Moore. Standardizing the next version of Ada. Ada User Journal, 25(2):91–??, June 2004. CODEN AU-JOET. ISSN 1381-6551.

Arkwright:2004:AAI

[353] Louise Arkwright. Ada academic initiative in Paris. Ada User Journal, 25(2): 97–??, June 2004. CODEN AUJOET. ISSN 1381-6551.

Amey:2004:PES

[354] Peter Amey and Adrian J. Hilton. Practical experiences of safety- and securitycritical technologies. Ada User Journal, 25(2):98–??, June 2004. CODEN AU-JOET. ISSN 1381-6551.

Alves:2004:NPG

[355] Mário Amado Alves. No pointers, great programs. Ada User Journal, 25(2):107– ??, June 2004. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2004:AESb

[356] Anonymous. Ada–Europe 2004 sponsors. Ada User Journal, 25(2):110–??, June 2004. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2004:EPAb

[357] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 25(3): 112–??, September 2004. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2004:Eb

[358] Anonymous. Editorial. Ada User Journal, 25(3):113–??, September 2004. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2004:Nc

[359] Anonymous. News. Ada User Journal, 25(3):115–??, September 2004. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2004:CCb

[360] Anonymous. Conference calendar. Ada User Journal, 25(3):144–??, September 2004. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2004:FEb

[361] Anonymous. Forthcoming events. Ada User Journal, 25(3):151–??, September 2004. CODEN AUJOET. ISSN 1381-6551.

Moore:2004:UAS

[362] James W. Moore. Update on Ada standardization. Ada User Journal, 25(3): 155–??, September 2004. CODEN AU-JOET. ISSN 1381-6551.

Rosen:2004:DWS

[363] Jean-Pierre Rosen. Developing a Web server in Ada with AWS. Ada User Journal, 25(3):157–??, September 2004. CO-DEN AUJOET. ISSN 1381-6551.

Heaney:2004:CCL

[364] Matthew J. Heaney. The Charles Container Library. Ada User Journal, 25(3): 166–??, September 2004. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2004:AESc

[365] Anonymous. Ada–Europe 2004 sponsors. Ada User Journal, 25(3):174–??, September 2004. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2004:EPAc

[366] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 25(4): 176–??, December 2004. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2004:Ec

[367] Anonymous. Editorial. Ada User Journal, 25(4):177–??, December 2004. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2004:Nd

[368] Anonymous. News. Ada User Journal, 25(4):179–??, December 2004. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2004:CCc

[369] Anonymous. Conference calendar. Ada User Journal, 25(4):209–??, December 2004. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2004:FEc

[370] Anonymous. Forthcoming events. Ada User Journal, 25(4):216–??, December 2004. CODEN AUJOET. ISSN 1381-6551.

Brosgol:2004:RTJ

[371] Benjamin M. Brosgol. Real-time Java for Ada programmers. Ada User Journal, 25(4):220–??, December 2004. CO-DEN AUJOET. ISSN 1381-6551.

Barnes:2004:RAI

[372] John Barnes. Rationale for Ada 2005: Introduction. Ada User Journal, 25(4): 228–??, December 2004. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2004:AESd

[373] Anonymous. Ada–Europe 2004 sponsors. Ada User Journal, 25(4):246–??, December 2004. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:EPAa

[374] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 26(1): 2–??, March 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:Ea

[375] Anonymous. Editorial. Ada User Journal, 26(1):3–??, March 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:Na

[376] Anonymous. News. Ada User Journal, 26(1):5–??, March 2005. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2005:CCa

[377] Anonymous. Conference calendar. Ada User Journal, 26(1):32–??, March 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:FEa

[378] Anonymous. Forthcoming events. Ada User Journal, 26(1):40–??, March 2005. CODEN AUJOET. ISSN 1381-6551.

Barnes:2005:RAO

[379] John Barnes. Rationale for Ada 2005: 1 object oriented model. Ada User Journal, 26(1):45–??, March 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:AESa

[380] Anonymous. Ada–Europe 2004 sponsors. Ada User Journal, 26(1):64–??, March 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:EPAb

[381] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 26(2): 66–??, June 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:Eb

[382] Anonymous. Editorial. Ada User Journal, 26(2):67–??, June 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:Nb

[383] Anonymous. News. Ada User Journal, 26(2):69–??, June 2005. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2005:CCb

[384] Anonymous. Conference calendar. Ada User Journal, 26(2):93–??, June 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:FEb

[385] Anonymous. Forthcoming events. Ada User Journal, 26(2):101–??, June 2005. CODEN AUJOET. ISSN 1381-6551.

Barnes:2005:RAA

[386] John Barnes. Rationale for Ada 2005: 2 access types. Ada User Journal, 26(2): 104–??, June 2005. CODEN AUJOET. ISSN 1381-6551.

Barnes:2005:RAS

[387] John Barnes. Rationale for Ada 2005: 3 structure and visibility. Ada User Journal, 26(2):120–??, June 2005. CODEN AUJOET. ISSN 1381-6551.

Ramachandran:2005:ARG

[388] Muthu Ramachandran. Ada reuse guidelines. Ada User Journal, 26(2): 135–??, June 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:AESb

[389] Anonymous. Ada–Europe 2005 sponsors. Ada User Journal, 26(2):144–??, June 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:EPAc

[390] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 26(3): 146–??, September 2005. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2005:Ec

[391] Anonymous. Editorial. Ada User Journal, 26(3):147–??, September 2005. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2005:Nc

[392] Anonymous. News. Ada User Journal, 26(3):149–??, September 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:CCc

[393] Anonymous. Conference calendar. Ada User Journal, 26(3):171–??, September 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:FEc

[394] Anonymous. Forthcoming events. Ada User Journal, 26(3):177–??, September 2005. CODEN AUJOET. ISSN 1381-6551.

Barnes:2005:RAT

[395] John Barnes. Rationale for Ada 2005: 4 tasking and real-time. Ada User Journal, 26(3):180–??, September 2005. CO-DEN AUJOET. ISSN 1381-6551.

Barnes:2005:RAE

[396] John Barnes. Rationale for Ada 2005: 5 exceptions, generics etc. Ada User Journal, 26(3):198–??, September 2005. CO-DEN AUJOET. ISSN 1381-6551.

Marriott:2005:ABF

[397] Alan Marriott and Urs Maurer. Ada bug finder. Ada User Journal, 26(3):214– ??, September 2005. CODEN AUJOET. ISSN 1381-6551.

Stadlmann:2005:ADB

[398] Burkhard Stadlmann. Ada development for a basic train control system for regional branch lines. Ada User Journal, 26(3):220–??, September 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:AESc

[399] Anonymous. Ada–Europe 2005 sponsors. Ada User Journal, 26(3):224–??, September 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:EPAd

[400] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 26(4): 226–??, December 2005. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2005:Ed

[401] Anonymous. Editorial. Ada User Journal, 26(4):227–??, December 2005. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2005:Nd

[402] Anonymous. News. Ada User Journal, 26(4):229–??, December 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:CCd

[403] Anonymous. Conference calendar. Ada User Journal, 26(4):253–??, December 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2005:FEd

[404] Anonymous. Forthcoming events. Ada User Journal, 26(4):260–??, December 2005. CODEN AUJOET. ISSN 1381-6551.

Barnes:2005:RAP

[405] John Barnes. Rationale for Ada 2005: 6 predefined library. Ada User Journal, 26(4):265–??, December 2005. CODEN AUJOET. ISSN 1381-6551.

Barnes:2005:RAC

[406] John Barnes. Rationale for Ada 2005: 6a containers. Ada User Journal, 26(4):

282–??, December 2005. CODEN AU-JOET. ISSN 1381-6551.

Sandberg:2005:LTS

[407] Per Sandberg and Rei Stråhle. Living in towers — the story of multi project system builds. Ada User Journal, 26(4): 306–??, December 2005. CODEN AU-JOET. ISSN 1381-6551.

Rosen:2005:BIS

[408] Jean-Pierre Rosen. On the benefits for industrials of sponsoring free software developments. Ada User Journal, 26(4): 308–??, December 2005. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2005:AESd

[409] Anonymous. Ada–Europe 2005 sponsors. Ada User Journal, 26(4):312–??, December 2005. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:EPAa

[410] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 27(1): 2–??, March 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:Ea

[411] Anonymous. Editorial. Ada User Journal, 27(1):3–??, March 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:Na

[412] Anonymous. News. Ada User Journal, 27(1):5–??, March 2006. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2006:CCa

[413] Anonymous. Conference calendar. Ada User Journal, 27(1):35–??, March 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:FEa

[414] Anonymous. Forthcoming events. Ada User Journal, 27(1):43–??, March 2006. CODEN AUJOET. ISSN 1381-6551.

Barnes:2006:RAE

[415] John Barnes. Rationale for Ada 2005: Epilogue. Ada User Journal, 27(1):46– ??, March 2006. CODEN AUJOET. ISSN 1381-6551.

Ramachandran:2006:GIA

[416] Muthu Ramachandran. Generating and improving Ada components for reuse. Ada User Journal, 27(1):56–??, March 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:AESa

[417] Anonymous. Ada–Europe 2005 sponsors. Ada User Journal, 27(1):64–??, March 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:EPAb

[418] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 27(2): 66–??, June 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:Eb

[419] Anonymous. Editorial. Ada User Journal, 27(2):67–??, June 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:Nb

[420] Anonymous. News. Ada User Journal, 27(2):69–??, June 2006. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2006:CCb

[421] Anonymous. Conference calendar. Ada User Journal, 27(2):97–??, June 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:FEb

[422] Anonymous. Forthcoming events. Ada User Journal, 27(2):106–??, June 2006. CODEN AUJOET. ISSN 1381-6551.

Varaprasad:2006:NSP

[423] G. Varaprasad, R. S. D. Wahidabanu, and P. Venkataram. A new strategy pattern for OO technology. Ada User Journal, 27(2):110–??, June 2006. CODEN AUJOET. ISSN 1381-6551.

Fairlamb:2006:ACU

[424] K Fairlamb. Ada Conference UK 2006.

Ada User Journal, 27(2):118–??, June 2006. CODEN AUJOET. ISSN 1381-6551.

Brandon:2006:AME

[425] A. S. Brandon. Ada market in 2005 entails at least a $5.6 billion investment. Ada User Journal, 27(2):123–??, June 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:AESb

[426] Anonymous. Ada–Europe 2006 sponsors. Ada User Journal, 27(2):128–??, June 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:EPAc

[427] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 27(3): 130–??, September 2006. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2006:Ec

[428] Anonymous. Editorial. Ada User Journal, 27(3):131–??, September 2006. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2006:Nc

[429] Anonymous. News. Ada User Journal, 27(3):133–??, September 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:CCc

[430] Anonymous. Conference calendar. Ada User Journal, 27(3):167–??, September 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:FEc

[431] Anonymous. Forthcoming events. Ada User Journal, 27(3):174–??, September 2006. CODEN AUJOET. ISSN 1381-6551.

Leroy:2006:MLD

[432] P Leroy. Memories of a language designer. Ada User Journal, 27(3):181– ??, September 2006. CODEN AUJOET. ISSN 1381-6551.

Kleidermacher:2006:DRS

[433] D. N. Kleidermacher. Developing reliable software rapidly. Ada User Journal, 27(3):184–??, September 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:AESc

[434] Anonymous. Ada–Europe 2006 sponsors. Ada User Journal, 27(3):192–??, September 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:EPAd

[435] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 27(4): 194–??, December 2006. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2006:Ed

[436] Anonymous. Editorial. Ada User Journal, 27(4):195–??, December 2006. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2006:Nd

[437] Anonymous. News. Ada User Journal, 27(4):197–??, December 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:CCd

[438] Anonymous. Conference calendar. Ada User Journal, 27(4):232–??, December 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:FEd

[439] Anonymous. Forthcoming events. Ada User Journal, 27(4):239–??, December 2006. CODEN AUJOET. ISSN 1381-6551.

Mahieux:2006:UCB

[440] J.-C. Mahieux, B. Maudry, and A. Foster. Using CORBA to bring new life to legacy Ada software: an experience report. Ada User Journal, 27(4):244– ??, December 2006. CODEN AUJOET. ISSN 1381-6551.

Klein:2006:PF

[441] J. Klein and D. Sotirovski. The publisher framework. Ada User Journal, 27(4):248–??, December 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2006:AESd

[442] Anonymous. Ada–Europe 2006 sponsors. Ada User Journal, 27(4):256–??, December 2006. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2007:EPAa

[443] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 28(1): 2–??, March 2007. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2007:Ea

[444] Anonymous. Editorial. Ada User Journal, 28(1):3–??, March 2007. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2007:Na

[445] Anonymous. News. Ada User Journal, 28(1):5–??, March 2007. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2007:CCa

[446] Anonymous. Conference calendar. Ada User Journal, 28(1):38–??, March 2007. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2007:FEa

[447] Anonymous. Forthcoming events. Ada User Journal, 28(1):45–??, March 2007. CODEN AUJOET. ISSN 1381-6551.

Comar:2007:ERB

[448] C. Comar and R. Berrendonner. ERB: a Ravenscar benchmarking framework. Ada User Journal, 28(1):53–??, March 2007. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2007:AESa

[449] Anonymous. Ada–Europe 2006 sponsors. Ada User Journal, 28(1):64–??, March 2007. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2007:EPAb

[450] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 28(2): 66–??, June 2007. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2007:Eb

[451] Anonymous. Editorial. Ada User Journal, 28(2):67–??, June 2007. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2007:Nb

[452] Anonymous. News. Ada User Journal, 28(2):69–??, June 2007. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2007:CCb

[453] Anonymous. Conference calendar. Ada User Journal, 28(2):102–??, June 2007. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2007:FEb

[454] Anonymous. Forthcoming events. Ada User Journal, 28(2):111–??, June 2007. CODEN AUJOET. ISSN 1381-6551.

delaPuente:2007:IRT

[455] J. A. de la Puente. 13th International Real-Time Ada Workshop (IRTAW-13). Ada User Journal, 28(2):117–??, June 2007. CODEN AUJOET. ISSN 1381-6551.

Glaister:2007:TDD

[456] W Glaister. Template driven design in UML and Ada 95. Ada User Journal, 28(2):122–??, June 2007. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2007:AESb

[457] Anonymous. Ada–Europe 2007 sponsors. Ada User Journal, 28(2):128–??, June 2007. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2007:EPAc

[458] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 28(3): 130–??, September 2007. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2007:Ec

[459] Anonymous. Editorial. Ada User Journal, 28(3):131–??, September 2007. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2007:Nc

[460] Anonymous. News. Ada User Journal, 28(3):133–??, September 2007. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2007:CCc

[461] Anonymous. Conference calendar. Ada User Journal, 28(3):166–??, September 2007. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2007:FEc

[462] Anonymous. Forthcoming events. Ada User Journal, 28(3):173–??, September 2007. CODEN AUJOET. ISSN 1381-6551.

Miranda:2007:TCO

[463] J. Miranda. Towards certification of object-oriented code with the GNAT compiler. Ada User Journal, 28(3):178– ??, September 2007. CODEN AUJOET. ISSN 1381-6551.

Black:2007:SES

[464] P. E. Black. SAMATE and evaluating static analysis tools. Ada User Journal, 28(3):184–??, September 2007. CODEN AUJOET. ISSN 1381-6551.

Bernat:2007:IOW

[465] G. Bernat, R. Davis, N. Merriam, J. Tuffen, A. Gardner, M. Bennett, and D. Armstrong. Identifying opportunities for worst-case execution time reduction in an avionics system. Ada User Journal, 28(3):189–??, September 2007. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2007:AEAa

[466] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 28(3):196– ??, September 2007. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2007:EPAd

[467] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 28(4): 198–??, December 2007. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2007:Ed

[468] Anonymous. Editorial. Ada User Journal, 28(4):199–??, December 2007. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2007:Nd

[469] Anonymous. News. Ada User Journal, 28(4):201–??, December 2007. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2007:CCd

[470] Anonymous. Conference calendar. Ada User Journal, 28(4):225–??, December 2007. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2007:FEd

[471] Anonymous. Forthcoming events. Ada User Journal, 28(4):232–??, December 2007. CODEN AUJOET. ISSN 1381-6551.

Furgel:2007:SSD

[472] I. Furgel and L. Hanke. Secure softwaredownload as part of a complex business process. Ada User Journal, 28(4):237– ??, December 2007. CODEN AUJOET. ISSN 1381-6551.

Vardanega:2007:SLI

[473] T. Vardanega and J. F. Ruiz. Session: Language issues. Ada User Journal, 28 (4):246–??, December 2007. CODEN AUJOET. ISSN 1381-6551.

Zerzelidis:2007:CEP

[474] A. Zerzelidis, A. Burns, and A. J. Wellings. Correcting the EDF protocol in Ada 2005. Ada User Journal, 28(4): 249–??, December 2007. CODEN AU-JOET. ISSN 1381-6551.

Wellings:2007:IOT

[475] A. J. Wellings and A. Burns. Integrating OOP and tasking — the missing requeue. Ada User Journal, 28(4):253– ??, December 2007. CODEN AUJOET. ISSN 1381-6551.

Uruena:2007:BHI

[476] S. Urueña and J. Zamorano. Building high-integrity distributed systems with Ravenscar restrictions. Ada User Journal, 28(4):259–??, December 2007. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2007:AG

[477] Anonymous. Ada gems. Ada User Journal, 28(4):267–??, December 2007. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2007:AEAb

[478] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 28(4):272–??, December 2007. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:EPAa

[479] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 29(1): 2–??, March 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:Ea

[480] Anonymous. Editorial. Ada User Journal, 29(1):3–??, March 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:Na

[481] Anonymous. News. Ada User Journal, 29(1):5–??, March 2008. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2008:CCa

[482] Anonymous. Conference calendar. Ada User Journal, 29(1):26–??, March 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:FEa

[483] Anonymous. Forthcoming events. Ada User Journal, 29(1):33–??, March 2008. CODEN AUJOET. ISSN 1381-6551.

Kleidermacher:2008:PAS

[484] D. Kleidermacher. Practical application of static analysis for embedded systems. Ada User Journal, 29(1):38–??, March 2008. CODEN AUJOET. ISSN 1381-6551.

Gonzalez-Harbour:2008:SPP

[485] M. González-Harbour and J. J.

Gutiérrez. Session: Programming patterns and libraries. Ada User Journal, 29(1):44–??, March 2008. CODEN AU-JOET. ISSN 1381-6551.

Wellings:2008:FRT

[486] A. J. Wellings and A. Burns. A framework for real-time utilities for Ada 2005. Ada User Journal, 29(1):47–??, March 2008. CODEN AUJOET. ISSN 1381-6551.

Burns:2008:PET

[487] A. Burns and A. J. Wellings. Programming execution-time servers in Ada 2005. Ada User Journal, 29(1):54–??, March 2008. CODEN AUJOET. ISSN 1381-6551.

Pulido:2008:ACP

[488] J. A. Pulido, J. A. de la Puente, J. Hugues, M. Bordin, and T. Vardanega. Ada 2005 code patterns for metamodel-based code generation. Ada User Journal, 29(1):59–??, March 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:AGa

[489] Anonymous. Ada gems. Ada User Journal, 29(1):66–??, March 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:AEAa

[490] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 29(1):72–??, March 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:EPAb

[491] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 29(2): 74–??, June 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:Eb

[492] Anonymous. Editorial. Ada User Journal, 29(2):75–??, June 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:Nb

[493] Anonymous. News. Ada User Journal, 29(2):77–??, June 2008. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2008:CCb

[494] Anonymous. Conference calendar. Ada User Journal, 29(2):98–??, June 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:FEb

[495] Anonymous. Forthcoming events. Ada User Journal, 29(2):106–??, June 2008. CODEN AUJOET. ISSN 1381-6551.

Burns:2008:SIE

[496] A. Burns and A. Wellings. Session: Implementation experience with Ada 2005. Ada User Journal, 29(2):112–??, June 2008. CODEN AUJOET. ISSN 1381-6551.

Uruena:2008:INA

[497] S. Urueña, J. Pulido, J. Redondo, and J. Zamorano. Implementing the new Ada 2005 real-time features on a bare board kernel. Ada User Journal, 29(2): 114–??, June 2008. CODEN AUJOET. ISSN 1381-6551.

Rivas:2008:OSS

[498] M. Aldea Rivas and M. González Harbour. Operating system support for execution time budgets for thread groups. Ada User Journal, 29(2):120–??, June 2008. CODEN AUJOET. ISSN 1381-6551.

Real:2008:SBA

[499] J. Real and S. Michell. Session: Beyond Ada 2005. Ada User Journal, 29(2):124– ??, June 2008. CODEN AUJOET. ISSN 1381-6551.

Wellings:2008:BAA

[500] A. Wellings and A. Burns. Beyond Ada 2005: Allocating tasks to processors in SMP systems. Ada User Journal, 29(2): 127–??, June 2008. CODEN AUJOET. ISSN 1381-6551.

Ward:2008:SSB

[501] M. Ward and N. C. Audsley. Suggestions for stream based parallel systems in Ada. Ada User Journal, 29(2):133– ??, June 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:AGb

[502] Anonymous. Ada gems. Ada User Journal, 29(2):140–??, June 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:AEAb

[503] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 29(2):144–??, June 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:EPAc

[504] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 29(3): 146–??, September 2008. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2008:Ec

[505] Anonymous. Editorial. Ada User Journal, 29(3):147–??, September 2008. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2008:Nc

[506] Anonymous. News. Ada User Journal, 29(3):149–??, September 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:CCc

[507] Anonymous. Conference calendar. Ada User Journal, 29(3):171–??, September 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:FEc

[508] Anonymous. Forthcoming events. Ada User Journal, 29(3):178–??, September 2008. CODEN AUJOET. ISSN 1381-

Brosgol:2008:SAO

[509] B. Brosgol and M. Aldea. Session: Ada and other standards. Ada User Journal, 29(3):186–??, September 2008. CODEN AUJOET. ISSN 1381-6551.

Michell:2008:IAO

[510] S. Michell. Interfacing Ada to operating systems. Ada User Journal, 29(3):188– ??, September 2008. CODEN AUJOET. ISSN 1381-6551.

delaPuente:2008:SCP

[511] J. A. de la Puente and S. Urueña. Session: Conclusions and plans for next IRTAW. Ada User Journal, 29(3):192– ??, September 2008. CODEN AUJOET. ISSN 1381-6551.

Real:2008:SRA

[512] J. Real and L. M. Pinho. Session report: Ada and software engineering education. Ada User Journal, 29(3):196– ??, September 2008. CODEN AUJOET. ISSN 1381-6551.

Schonberg:2008:PAS

[513] E. Schonberg and R. Dewar. A principled approach to software engineering education, or Java considered harmful. Ada User Journal, 29(3):200–??, September 2008. CODEN AUJOET. ISSN 1381-6551.

McCormick:2008:ASE

[514] J. W. McCormick. Ada and software engineering education: One professor’s experiences. Ada User Journal, 29(3):203– ??, September 2008. CODEN AUJOET. ISSN 1381-6551.

Rosen:2008:AEI

[515] J.-P. Rosen. Is Ada education important? Ada User Journal, 29(3):208– ??, September 2008. CODEN AUJOET. ISSN 1381-6551.

Brandon:2008:UAS

[516] C. Brandon. Use of Ada in a student CubeSat project. Ada User Journal, 29 (3):213–??, September 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:AGc

[517] Anonymous. Ada gems. Ada User Journal, 29(3):217–??, September 2008. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2008:AEAc

[518] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 29(3):220– ??, September 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:EPAd

[519] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 29(4): 222–??, December 2008. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2008:Ed

[520] Anonymous. Editorial. Ada User Journal, 29(4):223–??, December 2008. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2008:Nd

[521] Anonymous. News. Ada User Journal, 29(4):225–??, December 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:CCd

[522] Anonymous. Conference calendar. Ada User Journal, 29(4):251–??, December 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2008:FEd

[523] Anonymous. Forthcoming events. Ada User Journal, 29(4):260–??, December 2008. CODEN AUJOET. ISSN 1381-6551.

Ausden:2008:ACI

[524] H. Ausden. Ada–C++ interfacing in the ERAM system. Ada User Journal, 29(4):267–??, December 2008. CODEN AUJOET. ISSN 1381-6551.

Cousins:2008:PNC

[525] J. Cousins. Porting naval command & control systems to Ada 2005. Ada User Journal, 29(4):271–??, December 2008. CODEN AUJOET. ISSN 1381-6551.

Rosen:2008:CIC

[526] J.-P. Rosen. A comparison of industrial coding rules. Ada User Journal, 29(4): 277–??, December 2008. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2008:AGd

[527] Anonymous. Ada gems. Ada User Journal, 29(4):283–??, December 2008. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2008:AEAd

[528] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 29(4):288–??, December 2008. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:EPAa

[529] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 30(1): 2–??, March 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:Ea

[530] Anonymous. Editorial. Ada User Journal, 30(1):3–??, March 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:Na

[531] Anonymous. News. Ada User Journal, 30(1):5–??, March 2009. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2009:CCa

[532] Anonymous. Conference calendar. Ada User Journal, 30(1):30–??, March 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:FEa

[533] Anonymous. Forthcoming events. Ada User Journal, 30(1):37–??, March 2009. CODEN AUJOET. ISSN 1381-6551.

Barnes:2009:TYA

[534] J. Barnes. Thirty years of the Ada User Journal. Ada User Journal, 30(1):43–??, March 2009. CODEN AUJOET. ISSN 1381-6551.

Moore:2009:PRI

[535] J. W. Moore and J. Benito. Progress report: ISO/IEC 24772, programming language vulnerabilities. Ada User Journal, 30(1):46–??, March 2009. CODEN AUJOET. ISSN 1381-6551.

Moore:2009:DSM

[536] B. J. Moore. Distributed status monitoring and control using remote buffers and Ada 2005. Ada User Journal, 30(1): 49–??, March 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:AGa

[537] Anonymous. Ada gems. Ada User Journal, 30(1):61–??, March 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:AEAa

[538] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 30(1):64–??, March 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:EPAb

[539] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 30(2): 66–??, June 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:Eb

[540] Anonymous. Editorial. Ada User Journal, 30(2):67–??, June 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:Nb

[541] Anonymous. News. Ada User Journal, 30(2):69–??, June 2009. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2009:CCb

[542] Anonymous. Conference calendar. Ada User Journal, 30(2):98–??, June 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:FEb

[543] Anonymous. Forthcoming events. Ada User Journal, 30(2):106–??, June 2009. CODEN AUJOET. ISSN 1381-6551.

Fairlamb:2009:AUC

[544] K. Fairlamb. Ada UK Conference 2009.

Ada User Journal, 30(2):111–??, June 2009. CODEN AUJOET. ISSN 1381-6551.

Harbaugh:2009:PBR

[545] J. S. Harbaugh. Pattern-based refactoring shrinks maintenance costs. Ada User Journal, 30(2):115–??, June 2009. CO-DEN AUJOET. ISSN 1381-6551.

Ochem:2009:BCL

[546] Q. Ochem. Building cross language applications with Ada. Ada User Journal, 30(2):119–??, June 2009. CODEN AU-JOET. ISSN 1381-6551.

Davis:2009:ETA

[547] R. I. Davis and I. Broster. Execution time: Analysis, verification and optimization for reliable systems. Ada User Journal, 30(2):121–??, June 2009. CO-DEN AUJOET. ISSN 1381-6551.

Rogers:2009:SFT

[548] P. Rogers. Software fault tolerance. Ada User Journal, 30(2):125–??, June 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:AGb

[549] Anonymous. Ada gems. Ada User Journal, 30(2):129–??, June 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:AEAb

[550] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 30(2):132–??, June 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:EPAc

[551] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 30(3): 134–??, September 2009. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2009:Ec

[552] Anonymous. Editorial. Ada User Journal, 30(3):135–??, September 2009. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2009:Nc

[553] Anonymous. News. Ada User Journal, 30(3):137–??, September 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:CCc

[554] Anonymous. Conference calendar. Ada User Journal, 30(3):163–??, September 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:FEc

[555] Anonymous. Forthcoming events. Ada User Journal, 30(3):169–??, September 2009. CODEN AUJOET. ISSN 1381-6551.

Tokar:2009:AEW

[556] J. L. Tokar. Ada Europe 2009 — Workshop on Software Vulnerabilities. Ada User Journal, 30(3):175–??, September 2009. CODEN AUJOET. ISSN 1381-6551.

Rosen:2009:RPL

[557] J.-P. Rosen. On removing programming language bias from the vulnerabilities document. Ada User Journal, 30(3): 177–??, September 2009. CODEN AU-JOET. ISSN 1381-6551.

Michell:2009:APL

[558] S. Michell. Ada and programming language vulnerabilities. Ada User Journal, 30(3):180–??, September 2009. CODEN AUJOET. ISSN 1381-6551.

Vardanega:2009:ALS

[559] T. Vardanega. An argument for language subsetting. Ada User Journal, 30 (3):185–??, September 2009. CODEN AUJOET. ISSN 1381-6551.

Burns:2009:CV

[560] A. Burns and A. J. Wellings. Concurrency vulnerabilities. Ada User Journal, 30(3):187–??, September 2009. CODEN AUJOET. ISSN 1381-6551.

Chapman:2009:PCE

[561] R. C. Chapman. The pros and cons of enumerating programming language vulnerabilities. Ada User Journal, 30(3): 190–??, September 2009. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2009:AGc

[562] Anonymous. Ada gems. Ada User Journal, 30(3):194–??, September 2009. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2009:AEAc

[563] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 30(3):200– ??, September 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:EPAd

[564] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 30(4): 202–??, December 2009. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2009:Ed

[565] Anonymous. Editorial. Ada User Journal, 30(4):203–??, December 2009. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2009:Nd

[566] Anonymous. News. Ada User Journal, 30(4):205–??, December 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:CCd

[567] Anonymous. Conference calendar. Ada User Journal, 30(4):230–??, December 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:FEd

[568] Anonymous. Forthcoming events. Ada User Journal, 30(4):236–??, December 2009. CODEN AUJOET. ISSN 1381-6551.

Vardanega:2009:BRA

[569] T. Vardanega. Book review: Ada for Software Engineers, by Mordechai Ben-Ari. Ada User Journal, 30(4):241–??, December 2009. CODEN AUJOET. ISSN 1381-6551.

Sobczak:2009:EEA

[570] M. Sobczak. Experiences in evaluating Ada with a pilot project. Ada User Journal, 30(4):243–??, December 2009. CO-DEN AUJOET. ISSN 1381-6551.

Delange:2009:CIO

[571] J. Delange, J. Hugues, and L. Pautet.

Couverture: an innovative open framework for coverage analysis of safety critical applications. Ada User Journal, 30 (4):248–??, December 2009. CODEN AUJOET. ISSN 1381-6551.

Vergnaud:2009:GCB

[572] T. Vergnaud, F. Gilliers, and H. Balp.

Generating component-based AADL applications with MyCCM-HI and Ocarina. Ada User Journal, 30(4):256–??, December 2009. CODEN AUJOET. ISSN 1381-6551.

Rivas:2009:AUG

[573] M. Aldea Rivas and M. González-Harbour. Ada user guide on MaRTE OS. Ada User Journal, 30(4):264–??, December 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2009:AEAd

[574] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 30(4):272–??, December 2009. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:EPAa

[575] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 31(1): 2–??, March 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:Ea

[576] Anonymous. Editorial. Ada User Journal, 31(1):3–??, March 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:QNDa

[577] Anonymous. Quarterly news digest. Ada User Journal, 31(1):5–??, March 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:CCa

[578] Anonymous. Conference calendar. Ada User Journal, 31(1):29–??, March 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:FEa

[579] Anonymous. Forthcoming events. Ada User Journal, 31(1):36–??, March 2010. CODEN AUJOET. ISSN 1381-6551.

Tooby:2010:OWC

[580] B. Tooby. Opinion: The word ’coding’ considered harmful. Ada User Journal, 31(1):41–??, March 2010. CODEN AU-JOET. ISSN 1381-6551.

Schonberg:2010:OHG

[581] E. Schonberg. Origins and history of GNAT. Ada User Journal, 31(1):42–??, March 2010. CODEN AUJOET. ISSN 1381-6551.

Barnes:2010:WDK

[582] J. Barnes. We don’t know nothing. Ada User Journal, 31(1):45–??, March 2010. CODEN AUJOET. ISSN 1381-6551.

Dobbing:2010:RTP

[583] B. Dobbing. The Ravenscar Tasking Profile for high integrity real-time programs. Ada User Journal, 31(1):47–??, March 2010. CODEN AUJOET. ISSN 1381-6551.

Barnes:2010:SWC

[584] J. Barnes. The SPARK way to correctness is via abstraction. Ada User Journal, 31(1):55–??, March 2010. CODEN AUJOET. ISSN 1381-6551.

Taft:2010:OOP

[585] S. T. Taft. Object-oriented programming enhancements in Ada 200Y. Ada User Journal, 31(1):65–??, March 2010. CODEN AUJOET. ISSN 1381-6551.

Leroy:2010:MLD

[586] P. Leroy. Memories of a language designer. Ada User Journal, 31(1):74–??, March 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:AEAa

[587] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 31(1):76–??, March 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:EPAb

[588] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 31(2): 78–??, June 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:Eb

[589] Anonymous. Editorial. Ada User Journal, 31(2):79–??, June 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:QNDb

[590] Anonymous. Quarterly news digest. Ada User Journal, 31(2):81–??, June 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:CCb

[591] Anonymous. Conference calendar. Ada User Journal, 31(2):109–??, June 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:FEb

[592] Anonymous. Forthcoming events. Ada User Journal, 31(2):116–??, June 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:ASP

[593] Anonymous. Annual student programming contest the Ada way. Ada User Journal, 31(2):121–??, June 2010. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2010:WLC

[594] Anonymous. WG9 letter to the community: Maintenance and revision of the Ada programming language. Ada User Journal, 31(2):122–??, June 2010. CO-DEN AUJOET. ISSN 1381-6551.

Lopez:2010:EPD

[595] J. López, Á. Esquinas, J. Zamorano, and J. A. de la Puente. Experience in programming device drivers with the Ravenscar profile. Ada User Journal, 31(2):123–??, June 2010. CODEN AU-JOET. ISSN 1381-6551.

Baillon:2010:ERS

[596] C. Baillon and S. Bouchez-Mongardé. Executable requirements in a safetycritical context with Ada. Ada User Journal, 31(2):131–??, June 2010. CO-DEN AUJOET. ISSN 1381-6551.

Sobczak:2010:PCA

[597] M. Sobczak. Polymorphic callbacks for Ada/C++ bindings. Ada User Journal, 31(2):136–??, June 2010. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2010:AGa

[598] Anonymous. Ada gems. Ada User Journal, 31(2):140–??, June 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:AEAb

[599] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 31(2):148–??, June 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:EPAc

[600] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 31(3): 150–??, September 2010. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2010:Ec

[601] Anonymous. Editorial. Ada User Journal, 31(3):151–??, September 2010. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2010:QNDc

[602] Anonymous. Quarterly news digest. Ada User Journal, 31(3):153–??, September 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:CCc

[603] Anonymous. Conference calendar. Ada User Journal, 31(3):179–??, September 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:FEc

[604] Anonymous. Forthcoming events. Ada User Journal, 31(3):185–??, September 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:SPC

[605] Anonymous. Student programming contest “The Ada Way”. Ada User Journal, 31(3):189–??, September 2010. CODEN AUJOET. ISSN 1381-6551.

Burns:2010:ASVa

[606] A. Burns and J. L. Tokar. Ada and the Software Vulnerabilities Project. Ada User Journal, 31(3):191–??, September 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:AGb

[607] Anonymous. Ada gems. Ada User Journal, 31(3):217–??, September 2010. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2010:AEAc

[608] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 31(3):220– ??, September 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:AESa

[609] Anonymous. Ada–Europe 2010 sponsors. Ada User Journal, 31(3):c2–??, September 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:EPAd

[610] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 31(4): 222–??, December 2010. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2010:Ed

[611] Anonymous. Editorial. Ada User Journal, 31(4):223–??, December 2010. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2010:QNDd

[612] Anonymous. Quarterly news digest. Ada User Journal, 31(4):225–??, December 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:CCd

[613] Anonymous. Conference calendar. Ada User Journal, 31(4):250–??, December 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:FEd

[614] Anonymous. Forthcoming events. Ada User Journal, 31(4):257–??, December 2010. CODEN AUJOET. ISSN 1381-6551.

Burns:2010:MSS

[615] A. Burns and A. J. Wellings. Multiprocessor systems session summary. Ada User Journal, 31(4):263–??, December 2010. CODEN AUJOET. ISSN 1381-6551.

Vardanega:2010:SSL

[616] T. Vardanega, M. González-Harbour, and L. M. Pinho. Session summary: Language and distribution issues. Ada User Journal, 31(4):266–??, December 2010. CODEN AUJOET. ISSN 1381-6551.

Michell:2010:CIR

[617] S. Michell and J. Real. Conclusions of the 14th International Real-Time Ada Workshop. Ada User Journal, 31(4): 273–??, December 2010. CODEN AU-JOET. ISSN 1381-6551.

Burns:2010:PRI

[618] A. Burns. Progress report from the 14th International Real-Time Ada Workshop — IRTAW14. Ada User Journal, 31(4): 275–??, December 2010. CODEN AU-JOET. ISSN 1381-6551.

Burns:2010:ASVb

[619] A. Burns and J. L. Tokar. Ada and the Software Vulnerabilities Project: the SPARK Annex. Ada User Journal, 31 (4):278–??, December 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:AGc

[620] Anonymous. Ada gems. Ada User Journal, 31(4):278–??, December 2010. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2010:AEAd

[621] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 31(4):296–??, December 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2010:AESb

[622] Anonymous. Ada–Europe 2010 sponsors. Ada User Journal, 31(4):c2–??, December 2010. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:EPAa

[623] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 32(1): 2, March 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:Ea

[624] Anonymous. Editorial. Ada User Journal, 32(1):3–4, March 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:QNDa

[625] Anonymous. Quarterly news digest. Ada User Journal, 32(1):5–29, March 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:CCa

[626] Anonymous. Conference calendar. Ada User Journal, 32(1):30–36, March 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:SPC

[627] Anonymous. Student programming contest “The Ada Way”. Ada User Journal, 32(1):40–42, March 2011. CODEN AU-JOET. ISSN 1381-6551.

Moore:2011:CWS

[628] B. J. Moore. A comparison of worksharing, work-seeking, and work-stealing parallelism strategies using Paraffin with Ada 2005. Ada User Journal, 32(1): 43–48, March 2011. CODEN AUJOET. ISSN 1381-6551.

Taft:2011:DPP

[629] S. T. Taft. Designing ParaSail — Parallel Specification and Implementation Language. Ada User Journal, 32(1):49– 59, March 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:AGa

[630] Anonymous. Ada gems. Ada User Journal, 32(1):60–63, March 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:AEAa

[631] Anonymous. Ada–Europe Associate Members (National Ada Organizations). Ada User Journal, 32(1):64–??, March 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:AESa

[632] Anonymous. Ada–Europe 2010 sponsors. Ada User Journal, 32(1):C3, March 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:EPAb

[633] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 32(2): 66, June 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:Eb

[634] Anonymous. Editorial. Ada User Journal, 32(2):67–68, June 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:QNDb

[635] Anonymous. Quarterly news digest. Ada User Journal, 32(2):69–96, June 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:CCb

[636] Anonymous. Conference calendar. Ada User Journal, 32(2):97–102, June 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:FEa

[637] Anonymous. Forthcoming events. Ada User Journal, 32(2):103–106, June 2011. CODEN AUJOET. ISSN 1381-6551.

Bridges:2011:ISP

[638] R. Bridges, F. Dordowsky, and H. Tschöpe. Implementing a software product line for a complex avionics system in Ada 83. Ada User Journal, 32(2): 107–115, June 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:AGb

[639] Anonymous. Ada gems. Ada User Journal, 32(2):116–127, June 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:AEAb

[640] Anonymous. Ada–Europe Associate Members (National Ada Organizations). Ada User Journal, 32(1):128–??, March 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:AESb

[641] Anonymous. Ada–Europe 2011 sponsors. Ada User Journal, 32(1):C3, March 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:EPAc

[642] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 32 (3):130, September 2011. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2011:Ec

[643] Anonymous. Editorial. Ada User Journal, 32(3):131–132, September 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:QNDc

[644] Anonymous. Quarterly news digest. Ada User Journal, 32(3):133–153, September 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:CCc

[645] Anonymous. Conference calendar. Ada User Journal, 32(3):154–157, September 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:FEb

[646] Anonymous. Forthcoming events. Ada User Journal, 32(3):158–163, September 2011. CODEN AUJOET. ISSN 1381-

Barnes:2011:RAI

[647] J. Barnes. Rationale for Ada 2012: Introduction. Ada User Journal, 32(3): 164–181, September 2011. CODEN AU-JOET. ISSN 1381-6551.

Thornley:2011:IHI

[648] P. Thornley. The implementation of high integrity data structures. Ada User Journal, 32(3):182–187, September 2011. CODEN AUJOET. ISSN 1381-6551.

Andersen:2011:CUA

[649] J. S. Andersen. Crimeville — using Ada inside an on-line multi-user game. Ada User Journal, 32(3):188–193, September 2011. CODEN AUJOET. ISSN 1381-6551.

Bradley:2011:AUG

[650] P. J. Bradley, J. A. de la Puente, and J. Zamorano. Ada user guide for LEGO MINDSTORMS NXT. Ada User Journal, 32(3):194–203, September 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:AEAc

[651] Anonymous. Ada–Europe Associate Members (National Ada Organizations). Ada User Journal, 32(3):204– ??, September 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:AESc

[652] Anonymous. Ada–Europe 2011 Sponsors. Ada User Journal, 32(3):C3, September 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:EPAd

[653] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 32 (4):206, December 2011. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2011:Ed

[654] Anonymous. Editorial. Ada User Journal, 32(4):207–208, December 2011. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2011:QNDd

[655] Anonymous. Quarterly news digest. Ada User Journal, 32(4):209–234, December 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:CCd

[656] Anonymous. Conference calendar. Ada User Journal, 32(4):235–240, December 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:FEc

[657] Anonymous. Forthcoming events. Ada User Journal, 32(4):241–246, December 2011. CODEN AUJOET. ISSN 1381-6551.

Barnes:2011:RAC

[658] J. Barnes. Rationale for Ada 2012: 1 Contracts and aspects. Ada User Journal, 32(4):247–264, December 2011. CO-DEN AUJOET. ISSN 1381-6551.

Gregertsen:2011:IUN

[659] K. N. Gregertsen and A. Skavhaug. Implementation and usage of the new Ada 2012 execution time control features. Ada User Journal, 32(4):265–275, December 2011. CODEN AUJOET. ISSN 1381-6551.

Rivas:2011:IRT

[660] M. Aldea Rivas. 15th International Real-Time Ada Workshop (IRTAW-15). Ada User Journal, 32(4):276–280, December 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:AGc

[661] Anonymous. Ada gems. Ada User Journal, 32(4):281–283, December 2011. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2011:AEAd

[662] Anonymous. Ada–Europe Associate Members (National Ada Organizations). Ada User Journal, 32(4):284–??, December 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2011:AESd

[663] Anonymous. Ada–Europe 2011 sponsors. Ada User Journal, 32(4):C3, December 2011. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:EPAa

[664] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 33(1): 2, March 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:Ea

[665] Anonymous. Editorial. Ada User Journal, 33(1):3–4, March 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:QNDa

[666] Anonymous. Quarterly news digest. Ada User Journal, 33(1):5–31, March 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:CCa

[667] Anonymous. Conference calendar. Ada User Journal, 33(1):32–39, March 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:FEa

[668] Anonymous. Forthcoming events. Ada User Journal, 33(1):40–44, March 2012. CODEN AUJOET. ISSN 1381-6551.

Barnes:2012:RAE

[669] J. G. P. Barnes. Rationale for Ada 2012:

2 Expressions. Ada User Journal, 33(1): 45–53, March 2012. CODEN AUJOET. ISSN 1381-6551.

Sanden:2012:ELM

[670] B. Sandén. Entity-life modeling: Designing reactive software architectures to the strengths of tasks. Ada User Journal, 33(1):54–62, March 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:AGa

[671] Anonymous. Ada gems. Ada User Journal, 33(1):63–67, March 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:AEAa

[672] Anonymous. Ada–Europe Associate Members (National Ada Organizations). Ada User Journal, 33(1):68–??, March 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:AESa

[673] Anonymous. Ada–Europe 2011 Sponsors. Ada User Journal, 33(1):C3, March 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:EPAb

[674] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 33(2): 70, June 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:Eb

[675] Anonymous. Editorial. Ada User Journal, 33(2):71–72, June 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:QNDb

[676] Anonymous. Quarterly news digest. Ada User Journal, 33(2):73–93, June 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:CCb

[677] Anonymous. Conference calendar. Ada User Journal, 33(2):94–98, June 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:FEb

[678] Anonymous. Forthcoming events. Ada User Journal, 33(2):99–102, June 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:PRA

[679] Anonymous. Press release: Ada 2012 language standard submitted to ISO. Ada User Journal, 33(2):103–105, June 2012. CODEN AUJOET. ISSN 1381-6551.

Barnes:2012:RAS

[680] J. G. P. Barnes. Rationale for Ada 2012: 3 Structure and visibility. Ada User Journal, 33(2):106–114, June 2012. CO-DEN AUJOET. ISSN 1381-6551.

Palm:2012:UMD

[681] S. Palm. Use of model driven code generation on the ASIM Project. Ada User Journal, 33(2):115–123, June 2012. CO-DEN AUJOET. ISSN 1381-6551.

Jennings:2012:BUS

[682] T. J. Jennings. The benefits of using SPARK for high-assurance software. Ada User Journal, 33(2):124–126, June 2012. CODEN AUJOET. ISSN 1381-6551.

Jennings:2012:UPG

[683] T. J. Jennings. The use of proof and generics in SPARK. Ada User Journal, 33(2):127–131, June 2012. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2012:AGb

[684] Anonymous. Ada gems. Ada User Journal, 33(2):132–135, June 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:AEAb

[685] Anonymous. Ada–Europe Associate Members (National Ada Organizations). Ada User Journal, 33(2):136–??, June 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:AESb

[686] Anonymous. Ada–Europe 2012 sponsors. Ada User Journal, 33(2):C3, June 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:EPAc

[687] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 33 (3):138, September 2012. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2012:Ec

[688] Anonymous. Editorial. Ada User Journal, 33(3):139–140, September 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:QNDc

[689] Anonymous. Quarterly news digest. Ada User Journal, 33(3):141–164, September 2012. CODEN AUJOET. ISSN 1381-

Anonymous:2012:CCc

[690] Anonymous. Conference calendar. Ada User Journal, 33(3):165–169, September 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:FEc

[691] Anonymous. Forthcoming events. Ada User Journal, 33(3):170–174, September 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:PRF

[692] Anonymous. Press release: First ‘Ada Way’ award winners and ‘Try and Beat Me’ Challenge. Ada User Journal, 33 (3):175–177, September 2012. CODEN AUJOET. ISSN 1381-6551.

Barnes:2012:RAT

[693] J. G. P. Barnes. Rationale for Ada 2012: 4 Tasking and real-time. Ada User Journal, 33(3):178–186, September 2012. CODEN AUJOET. ISSN 1381-6551.

Vardanega:2012:WLT

[694] T. Vardanega, F. Gasperoni, E. Plödereder, J. M. Martínez Rodríguez, B. Meyer, and A. Llemosí. What is language technology in our time? Ada User Journal, 33(3):187–203, September 2012. CO-DEN AUJOET. ISSN 1381-6551.

Bundgaard:2012:RSP

[695] J. Bundgaard, A. Rodríguez, S. Palm, R. E. Sward, and J. Ruiz. Reliable software, a perspective from industry. Ada User Journal, 33(3):204–211, September 2012. CODEN AUJOET. ISSN 1381-6551.

Reyes:2012:AWD

[696] R. Aguirre Reyes, A. Graziano, M. Teoli, and A. Zuccato. The Ada Way: development of a soccer simulator. Ada User Journal, 33(3):212–216, September 2012. CODEN AUJOET. ISSN 1381-6551.

Dordowsky:2012:CCG

[697] F. Dordowsky, R. Bridges, and H. Tschöpe. Combining code generation and Ada generics to implement a software product line. Ada User Journal, 33 (3):217–226, September 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:AGc

[698] Anonymous. Ada gems. Ada User Journal, 33(3):227–229, September 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:AEAc

[699] Anonymous. Ada–Europe Associate Members (National Ada Organizations). Ada User Journal, 33(3):230– ??, September 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:AESc

[700] Anonymous. Ada–Europe 2012 sponsors. Ada User Journal, 33(3):C3, September 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:EPB

[701] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 33 (4):232, December 2012. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2012:E

[702] Anonymous. Editorial. Ada User Journal, 33(4):233–234, December 2012. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2012:QND

[703] Anonymous. Quarterly news digest. Ada User Journal, 33(4):235–??, December 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:CC

[704] Anonymous. Conference calendar. Ada User Journal, 33(4):259–??, December 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:FE

[705] Anonymous. Forthcoming events. Ada User Journal, 33(4):264–??, December 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:ATB

[706] Anonymous. AdaWay try-and-beat-me challenge. Ada User Journal, 33(4):270– ??, December 2012. CODEN AUJOET. ISSN 1381-6551.

Barnes:2012:RAI

[707] J. G. P. Barnes. Rationale for Ada 2012: 5 iterators, pools, etc. Ada User Journal, 33(4):271–??, December 2012. CODEN AUJOET. ISSN 1381-6551.

Kamei:2012:TSV

[708] F. Kamei, D. C. Carta, I. Broster, and W. Lunniss. Tool support for verification of software timing and stack usage for a DO-178B Level A system. Ada User Journal, 33(4):289–??, December 2012. CODEN AUJOET. ISSN 1381-6551.

Ferrero:2012:IHS

[709] F. Ferrero, E. Alaña, A. I. Rodríguez, J. Zamorano, and J. A. de la Puente. Including hardware/software co-design in the ASSERT model driven engineering process. Ada User Journal, 33(4):293– ??, December 2012. CODEN AUJOET. ISSN 1381-6551.

Rego:2012:AUG

[710] P. V. Rego. Ada user guide: Integrating 8-bit AVR micro-controllers in Ada. Ada User Journal, 33(4):301–??, December 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2012:AEA

[711] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 33(4):306, December 2012. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:EPBa

[712] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 34(1): 2, March 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:Ea

[713] Anonymous. Editorial. Ada User Journal, 34(1):3–4, March 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:QNDa

[714] Anonymous. Quarterly news digest. Ada User Journal, 34(1):5–??, March 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:CCa

[715] Anonymous. Conference calendar. Ada User Journal, 34(1):22–??, March 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:FEa

[716] Anonymous. Forthcoming events. Ada User Journal, 34(1):29–??, March 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:ALS

[717] Anonymous. Ada 2012 language standard approved by ISO. Ada User Journal, 34(1):35–??, March 2013. CODEN AUJOET. ISSN 1381-6551.

Barnes:2013:RAP

[718] J. G. P. Barnes. Rationale for Ada 2012: 6 predefined library. Ada User Journal, 34(1):38–??, March 2013. CODEN AU-JOET. ISSN 1381-6551.

Rosen:2013:DCC

[719] J.-P. Rosen. Designing and checking coding standards for Ada. Ada User Journal, 34(1):45–??, March 2013. CO-DEN AUJOET. ISSN 1381-6551.

Rivas:2013:AAS

[720] M. Aldea Rivas. Advanced Ada support for real-time programming. Ada User Journal, 34(1):49–??, March 2013. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2013:AGa

[721] Anonymous. Ada gems. Ada User Journal, 34(1):57–??, March 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:AEAa

[722] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 34(1), March 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:EPBb

[723] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 34(2): 62, June 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:Eb

[724] Anonymous. Editorial. Ada User Journal, 34(2):63, June 2013. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2013:QNDb

[725] Anonymous. Quarterly news digest. Ada User Journal, 34(2):64–78, June 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:FEb

[726] Anonymous. Forthcoming events. Ada User Journal, 34(2):85–89, June 2013. CODEN AUJOET. ISSN 1381-6551.

Barnes:2013:RAC

[727] J. G. P. Barnes. Rationale for Ada 2012:

6a containers. Ada User Journal, 34(2): 90–107, June 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:OIR

[728] Anonymous. Overview of the 15th International Real-Time Ada Workshop (IRTAW 2011). Ada User Journal, 34 (2):108, June 2013. CODEN AUJOET. ISSN 1381-6551.

Real:2013:SSM

[729] J. Real and J. F. Ruiz. Session summary:

Multiprocessor issues, part 1. Ada User Journal, 34(2):109–111, June 2013. CO-DEN AUJOET. ISSN 1381-6551.

Wellings:2013:SSM

[730] A. Wellings and L. M. Pinho. Session summary: Multiprocessor issues, part 2 (resource control protocols). Ada User Journal, 34(2):112–116, June 2013. CO-DEN AUJOET. ISSN 1381-6551.

Burns:2013:SSLa

[731] A. Burns and T. Vardanega. Session summary: Language profile and application frameworks. Ada User Journal, 34(2):117–119, June 2013. CODEN AU-JOET. ISSN 1381-6551.

delaPuente:2013:SSC

[732] J. A. de la Puente and S. Michell. Session summary: Concurrency issues. Ada User Journal, 34(2):120–124, June 2013. CODEN AUJOET. ISSN 1381-6551.

Robinson:2013:RAC

[733] S. Robinson. Report: Ada Conference UK 2013. Ada User Journal, 34(2): 125–126, June 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:AGb

[734] Anonymous. Ada gems. Ada User Journal, 34(2):127–131, June 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:AEAb

[735] Anonymous. Ada–Europe Associate Members (National Ada Organizations). Ada User Journal, 34(2):132–??, June 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:AESa

[736] Anonymous. Ada–Europe 2013 sponsors. Ada User Journal, 34(2):c3, June 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:EPBc

[737] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 34 (3):134, September 2013. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2013:Ec

[738] Anonymous. Editorial. Ada User Journal, 34(3):135, September 2013. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2013:QNDc

[739] Anonymous. Quarterly news digest. Ada User Journal, 34(3):136–156, September 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:CCb

[740] Anonymous. Conference calendar. Ada User Journal, 34(3):157–161, September 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:FEc

[741] Anonymous. Forthcoming events. Ada User Journal, 34(3):162–166, September 2013. CODEN AUJOET. ISSN 1381-6551.

Barnes:2013:RAE

[742] J. G. P. Barnes. Rationale for Ada 2012: Epilogue. Ada User Journal, 34 (3):167–174, September 2013. CODEN AUJOET. ISSN 1381-6551.

Bigelow:2013:UGE

[743] D. Bigelow. Using the GNAT environment to maintain a large codebase inherited from another compilation system. Ada User Journal, 34(3):175–182, September 2013. CODEN AUJOET. ISSN 1381-6551.

Burns:2013:RIR

[744] A. Burns. Report: 16th International Real-Time Ada Workshop. Ada User Journal, 34(3):183–186, September 2013. CODEN AUJOET. ISSN 1381-6551.

Plodereder:2013:HUH

[745] E. Plödereder and J. Bundgaard. How to use the heap in real-time systems: Panel report. Ada User Journal, 34(3):

187–189, September 2013. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2013:AGc

[746] Anonymous. Ada gems. Ada User Journal, 34(3):190–191, September 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:AEAc

[747] Anonymous. Ada–Europe Associate Members (National Ada Organizations). Ada User Journal, 34(3):192– ??, September 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:AESb

[748] Anonymous. Ada–Europe 2013 sponsors. Ada User Journal, 34(3):c3, September 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:EPBd

[749] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 34 (4):194, December 2013. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2013:Ed

[750] Anonymous. Editorial. Ada User Journal, 34(4):195, December 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:LPA

[751] Anonymous. Letter from the President of Ada–Europe. Ada User Journal, 34 (4):196, December 2013. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2013:QNDd

[752] Anonymous. Quarterly news digest. Ada User Journal, 34(4):197–210, December 2013. CODEN AUJOET. ISSN 1381-

Anonymous:2013:CCc

[753] Anonymous. Conference calendar. Ada User Journal, 34(4):211–216, December 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:FEd

[754] Anonymous. Forthcoming events. Ada User Journal, 34(4):217–222, December 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:PRA

[755] Anonymous. Press release: Ada 2012 Language Rationale published. Ada User Journal, 34(4):223–225, December 2013. CODEN AUJOET. ISSN 1381-6551.

Andersen:2013:AA

[756] Jacob Sparre Andersen. Alice in Adaland. Ada User Journal, 34(4):223–225, December 2013. CODEN AUJOET. ISSN 1381-6551.

Pinho:2013:SSP

[757] L. M. Pinho, S. Michell, and B. Moore.

Session summary: Parallel and multicore systems. Ada User Journal, 34 (4):230, December 2013. CODEN AU-JOET. ISSN 1381-6551.

Burns:2013:SSLb

[758] A. Burns and A. Wellings. Session summary: Locking protocols. Ada User Journal, 34(4):237–238, December 2013. CODEN AUJOET. ISSN 1381-6551.

Vardanega:2013:SSI

[759] T. Vardanega and R. White. Session summary: Improvements to Ada. Ada User Journal, 34(4):239–241, December 2013. CODEN AUJOET. ISSN 1381-

Real:2013:SSO

[760] J. Real and J. A. de la Puente. Session summary: Open issues. Ada User Journal, 34(4):242, December 2013. CODEN AUJOET. ISSN 1381-6551.

Moy:2013:SR

[761] Y. Moy. SPARK 2014 rationale. Ada User Journal, 34(4):243–252, December 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:AGd

[762] Anonymous. Ada gems. Ada User Journal, 34(4):243–252, December 2013. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2013:AEAd

[763] Anonymous. Ada–Europe Associate Members (National Ada Organizations). Ada User Journal, 34(4):256–??, December 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2013:AESc

[764] Anonymous. Ada–Europe 2013 sponsors. Ada User Journal, 34(4):c3, December 2013. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:EPAa

[765] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 35(1): 2, March 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:Ea

[766] Anonymous. Editorial. Ada User Journal, 35(1):3, March 2014. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2014:QNDa

[767] Anonymous. Quarterly news digest. Ada User Journal, 35(1):4–??, March 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:CCa

[768] Anonymous. Conference calendar. Ada User Journal, 35(1):26–??, March 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:FEa

[769] Anonymous. Forthcoming events. Ada User Journal, 35(1):33–??, March 2014. CODEN AUJOET. ISSN 1381-6551.

Sargsyan:2014:RSB

[770] K. Sargsyan. Reliable software in bioinformatics: Sequence alignment with Coq, Ada and SPARK. Ada User Journal, 35(1):38–??, March 2014. CODEN AUJOET. ISSN 1381-6551.

Grein:2014:PUG

[771] C. K. W. Grein. Physical units with GNAT. Ada User Journal, 35(1):42–??, March 2014. CODEN AUJOET. ISSN 1381-6551.

Ekman:2014:TQS

[772] M. Ekman, H. Thane, D. Sundmark, and S. Larsson. Tool qualification for safety related systems. Ada User Journal, 35(1):47–??, March 2014. CODEN AUJOET. ISSN 1381-6551.

delaPuente:2014:ESB

[773] J. A. de la Puente, A. Alonso, J. Zamorano, J. Garrido, E. Salazar, and M. A. de Miguel. Experience in spacecraft on-board software development. Ada User Journal, 35(1):55–??, March 2014. CODEN AUJOET. ISSN 1381-6551.

Dross:2014:SRF

[774] C. Dross. SPARK 2014 rationale: Formal containers. Ada User Journal, 35 (1):61–??, March 2014. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2014:AG

[775] Anonymous. Ada gems. Ada User Journal, 35(1):65–??, March 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:AEAa

[776] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 35(1):68–??, March 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:AESa

[777] Anonymous. Ada–Europe 2013 sponsors. Ada User Journal, 35(1):ibc, March 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:EPAb

[778] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 35(2): 70, June 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:Eb

[779] Anonymous. Editorial. Ada User Journal, 35(2):71, June 2014. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2014:QNDb

[780] Anonymous. Quarterly news digest. Ada User Journal, 35(2):72–??, June 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:CCb

[781] Anonymous. Conference calendar. Ada User Journal, 35(2):94–??, June 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:FEb

[782] Anonymous. Forthcoming events. Ada User Journal, 35(2):100–??, June 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:PRA

[783] Anonymous. Press release: Ada 2012 Language Standard published in Springer’s LNCS and as free eBook. Ada User Journal, 35(2):104–??, June 2014. CODEN AUJOET. ISSN 1381-6551.

Mefteh:2014:FME

[784] M. Mefteh, N. Bouassida, and H. Ben-Abdallah. Feature model extraction from documented UML use case diagrams. Ada User Journal, 35(2):107–??, June 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:PWM

[785] Anonymous. Proceedings of the “Workshop on Mixed Criticality for Industrial Systems” of Ada–Europe 2014. Ada User Journal, 35(2):117–??, June 2014. CODEN AUJOET. ISSN 1381-6551.

Davis:2014:PPA

[786] R. Davis et al. PROXIMA: A probabilistic approach to the timing behaviour of mixed-criticality systems. Ada User Journal, 35(2):118–??, June 2014. CO-DEN AUJOET. ISSN 1381-6551.

Alonso:2014:TMC

[787] A. Alonso and E. Salazar. Toolset for mixed-criticality partitioned systems: Partitioning algorithm and extensibility support. Ada User Journal, 35 (2):123–??, June 2014. CODEN AU-JOET. ISSN 1381-6551.

Lindgren:2014:RLS

[788] P. Lindgren, D. Pereira, J. Eriksson, M. Linder, and L. M. Pinho. RTFM-lang static semantics for systems with mixed criticality. Ada User Journal, 35(2):128– ??, June 2014. CODEN AUJOET. ISSN 1381-6551.

Jan:2014:HCM

[789] M. Jan, L. Zaourar, V. Legout, and L. Pautet. Handling criticality mode change in time-triggered systems through linear programming. Ada User Journal, 35(2):133–??, June 2014. CO-DEN AUJOET. ISSN 1381-6551.

Cros:2014:MCS

[790] O. Cros, F. Fauberteau, L. George, and X. Li. Mixed criticality over switched Ethernet networks. Ada User Journal, 35(2):138–??, June 2014. CODEN AU-JOET. ISSN 1381-6551.

Cohen:2014:MCR

[791] A. Cohen, V. Perrelle, D. Potop-Butucaru, E. Soubiran, and Z. Zhang. Mixed criticality in railway systems: A case study on signaling application. Ada User Journal, 35(2):144–??, June 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:AEAb

[792] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 35(2):148–??, June 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:AESb

[793] Anonymous. Ada–europe sponsors. Ada User Journal, 35(2):ibc, June 2014. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2014:EPAc

[794] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 35 (3):150, September 2014. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2014:Ec

[795] Anonymous. Editorial. Ada User Journal, 35(3):151, September 2014. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2014:QNDc

[796] Anonymous. Quarterly news digest. Ada User Journal, 35(3):152–??, September 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:CCc

[797] Anonymous. Conference calendar. Ada User Journal, 35(3):172–??, September 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:FEc

[798] Anonymous. Forthcoming events. Ada User Journal, 35(3):178–??, September 2014. CODEN AUJOET. ISSN 1381-6551.

Rodriguez:2014:CSF

[799] A. Rodríguez. Critical software for the first European rail traffic management system. Ada User Journal, 35(3):186– ??, September 2014. CODEN AUJOET. ISSN 1381-6551.

Andersen:2014:PLJ

[800] J. S. Andersen. Privacy leaks in Java classes. Ada User Journal, 35(3):191– ??, September 2014. CODEN AUJOET. ISSN 1381-6551.

Pedro:2014:ITT

[801] A. M. Pedro. Implementation of task types in AVR-Ada. Ada User Journal, 35(3):194–??, September 2014. CODEN AUJOET. ISSN 1381-6551.

Efstathopoulos:2014:SRD

[802] P. Efstathopoulos. SPARK 2014 rationale: Data dependencies and information flow. Ada User Journal, 35(3):204– ??, September 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:AEAc

[803] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 35(3):208– ??, September 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:AESc

[804] Anonymous. Ada–Europe sponsors. Ada User Journal, 35(3):ibc, September 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:EPAd

[805] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 35 (4):210, December 2014. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2014:E

[806] Anonymous. Editorial. Ada User Journal, 35(4):211, December 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:QND

[807] Anonymous. Quarterly news digest. Ada User Journal, 35(4):212–233, December 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:CC

[808] Anonymous. Conference calendar. Ada User Journal, 35(4):234–239, December 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:FE

[809] Anonymous. Forthcoming events. Ada User Journal, 35(4):240–245, December 2014. CODEN AUJOET. ISSN 1381-6551.

Cholay:2014:ABD

[810] R. Cholay. AdDoc (beyond a document generator). Ada User Journal, 35 (4):246–248, December 2014. CODEN AUJOET. ISSN 1381-6551. Presented at the Industrial Track of Ada–Europe 2014.

Anonymous:2014:PWC

[811] Anonymous. Proceedings of the “Workshop on Challenges and New Approaches for Dependable and Cyber-Physical System Engineering” of Ada–Europe 2014. Ada User Journal, 35(4):249, December 2014. CODEN AUJOET. ISSN 1381-6551.

Cancila:2014:E

[812] D. Cancila and J-L. Gerstenmayer. Editorial. Ada User Journal, 35(4):250, December 2014. CODEN AUJOET. ISSN 1381-6551.

Attwood:2014:UCV

[813] K. Attwood, P. Conmy, and T. Kelly.

The use of controlled vocabularies and structured expressions in the assurance of CPS. Ada User Journal, 35(4):251– 258, December 2014. CODEN AU-JOET. ISSN 1381-6551.

David:2014:DRT

[814] V. David, A. Barbot, and D. Chabrol. Dependable real-time system and mixed criticality: Seeking safety, flexibility and efficiency with Kron-OS. Ada User Journal, 35(4):259–265, December 2014. CO-DEN AUJOET. ISSN 1381-6551.

Nakajima:2014:BCE

[815] S. Nakajima and M. Toyoshima. Behavioral contracts for energy consumption. Ada User Journal, 35(4):266–271, December 2014. CODEN AUJOET. ISSN 1381-6551.

Cancila:2014:FSU

[816] D. Cancila, E. Soubiran, and R. Passerone. Feasibility study in the use of contractbased approaches to deal with safetyrelated properties in CPS. Ada User Journal, 35(4):272–277, December 2014. CODEN AUJOET. ISSN 1381-6551.

Cancila:2014:RT

[817] D. Cancila, J-L. Gerstenmayer, C. Robinson, and L. Rioux. Round table. Ada User Journal, 35(4):278–279, December 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:AEA

[818] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 35(4):280–??, December 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2014:AES

[819] Anonymous. Ada–Europe sponsors. Ada User Journal, 35(4):ibc, December 2014. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2015:EPAa

[820] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 36(1): 2, March 2015. CODEN AUJOET. ISSN 1381-6551.

Pinho:2015:Ea

[821] Luís Miguel Pinho. Editorial. Ada User Journal, 36(1):3, March 2015. CODEN AUJOET. ISSN 1381-6551.

Andersen:2015:QNDa

[822] Jacob Sparre Andersen. Quarterly news digest. Ada User Journal, 36(1):4–23, March 2015. CODEN AUJOET. ISSN 1381-6551.

Craeynest:2015:CCa

[823] Dirk Craeynest. Conference calendar.

Ada User Journal, 36(1):24–31, March 2015. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2015:FEa

[824] Anonymous. Forthcoming events. Ada User Journal, 36(1):32–35, March 2015. CODEN AUJOET. ISSN 1381-6551.

Charman-Anderson:2015:ALV

[825] Suw Charman-Anderson. Ada Lovelace:

Victorian computing visionary. Ada User Journal, 36(1):35–41, March 2015. CODEN AUJOET. ISSN 1381-6551. Bicentennial Ada Lovelace articles.

Andersen:2015:PCA

[826] Jacob Sparre Andersen. Persistent containers with Ada 2012. Ada User Journal, 36(1):43–48, March 2015. CODEN AUJOET. ISSN 1381-6551. Presented at The Advances on Methods Special Session of Ada–Europe 2015.

Sachez-Ledesma:2015:TBC

[827] Francisco Sáchez-Ledesma, Juan Pastor, Diego Alonso, and Bárbara Ávarez. A task-based concurrency scheme for executing component-based applications. Ada User Journal, 36(1):49–55, March 2015. CODEN AUJOET. ISSN 1381-6551. Presented at The Advances on Methods Special Session of Ada–Europe 2015.

Anonymous:2015:AEAa

[828] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 36(1):56–??, March 2015. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2015:AESa

[829] Anonymous. Ada–Europe sponsors. Ada User Journal, 36(1):ibc, March 2015. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2015:EPAb

[830] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 36(2): 58, June 2015. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2015:Eb

[831] Anonymous. Editorial. Ada User Journal, 36(2):59, June 2015. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2015:QNDb

[832] Anonymous. Quarterly news digest. Ada User Journal, 36(2):60–77, June 2015. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2015:CCb

[833] Anonymous. Conference calendar. Ada User Journal, 36(2):78–84, June 2015. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2015:FEb

[834] Anonymous. Forthcoming events. Ada User Journal, 36(2):85–88, June 2015. CODEN AUJOET. ISSN 1381-6551.

Fuegi:2015:LBC

[835] J. Fuegi and J. Francis. Lovelace & Babbage and the creation of the 1843 ’Notes’. Ada User Journal, 36(2):89–98, June 2015. CODEN AUJOET. ISSN 1381-6551. Bicentennial Ada Lovelace articles.

Martignano:2015:SCA

[836] M. Martignano, A. Jung, T. Lehmann, and C. Schmidt. Source code analysis of flight software using a SonarQube based code quality platform. Ada User Journal, 36(2):99–104, June 2015. CODEN AUJOET. ISSN 1381-6551. from the Industrial Track of Ada–Europe 2015.

Baird:2015:SRS

[837] S. Baird, C. Dross, Y. Moy, T. Taft, and F. Schanda. Support of Ravenscar in SPARK 2014. Ada User Journal, 36(2): 105–112, June 2015. CODEN AUJOET. ISSN 1381-6551.

Moy:2015:SRG

[838] Y. Moy. SPARK 2014 rationale: Ghost code, object oriented programming and functional update. Ada User Journal, 36(2):113–115, June 2015. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2015:AEAb

[839] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 36(2):116–??, June 2015. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2015:AESb

[840] Anonymous. Ada–Europe sponsors. Ada User Journal, 36(2):ibc, June 2015. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2015:EPAc

[841] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 36 (3):118, September 2015. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2015:Ec

[842] Anonymous. Editorial. Ada User Journal, 36(3):119, September 2015. CO-DEN AUJOET. ISSN 1381-6551.

Anonymous:2015:QNDc

[843] Anonymous. Quarterly news digest. Ada User Journal, 36(3):120–140, September 2015. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2015:CCc

[844] Anonymous. Conference calendar. Ada User Journal, 36(3):141–145, September 2015. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2015:FEc

[845] Anonymous. Forthcoming events. Ada User Journal, 36(3):146–151, September 2015. CODEN AUJOET. ISSN 1381-6551.

Lovelace:2015:NTS

[846] A. A. Lovelace. 1842 notes to the translation of the sketch of The Analytical Engine. Ada User Journal, 36(3):152– 181, September 2015. CODEN AU-JOET. ISSN 1381-6551. Bicentennial Ada Lovelace articles.

Law:2015:EWC

[847] S. Law, M. Bennett, S. Hutchesson, I. Ellis, G. Bernat, A. Colin, and A. Coombes. Effective worst-case execution time analysis of DO178C Level A software. Ada User Journal, 36(3): 182–186, September 2015. CODEN AU-JOET. ISSN 1381-6551. Article from the Advances on Methods session of Ada– Europe 2015.

Hawkins:2015:NWM

[848] R. Hawkins, I. Habli, and T. Kelly. The need for a weaving model in assurance case automation. Ada User Journal, 36 (3):187–191, September 2015. CODEN AUJOET. ISSN 1381-6551. From the Architecture Centric Virtual Integration Workshop of Ada–Europe 2015.

Feiler:2015:ALR

[849] P. H. Feiler. Architecture-led requirements and safety analysis of an aircraft survivability situational awareness system. Ada User Journal, 36(3):192–195, September 2015. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2015:AEAc

[850] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 36(3):196– ??, September 2015. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2015:AESc

[851] Anonymous. Ada–Europe sponsors.

Ada User Journal, 36(3):ibc, September 2015. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2015:EPAd

[852] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 36 (4):198, December 2015. CODEN AU-JOET. ISSN 1381-6551.

Pinho:2015:Eb

[853] Luís Miguel Pinho. Editorial. Ada User Journal, 36(4):199, December 2015. CO-DEN AUJOET. ISSN 1381-6551.

Andersen:2015:QNDv

[854] Jacob Sparre Andersen. Quarterly news digest. Ada User Journal, 36(4):200– 215, December 2015. CODEN AU-JOET. ISSN 1381-6551.

Craeynest:2015:CCd

[855] Dirk Craeynest. Conference calendar. Ada User Journal, 36(4):216–220, December 2015. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2015:FEd

[856] Anonymous. Forthcoming events. Ada User Journal, 36(4):221–225, December 2015. CODEN AUJOET. ISSN 1381-6551.

RinoNesin:2015:ALS

[857] Gabriela Asli Rino Nesin. The Ada Lovelace symposium in Oxford. Ada User Journal, 36(4):225–230, December 2015. CODEN AUJOET. ISSN 1381-6551. Bicentennial Ada Lovelace article.

Barnes:2015:BAL

[858] John G. P. Barnes. From Byron to the Ada language. Ada User Journal, 36(4): 231–235, December 2015. CODEN AU-JOET. ISSN 1381-6551. Bicentennial Ada Lovelace article.

Gallina:2015:DRP

[859] Barbara Gallina and Luciana Provenzano. Deriving reusable process-based arguments from process models in the context of railway safety standards. Ada User Journal, 36(4):237–241, December 2015. CODEN AUJOET. ISSN 1381-6551. Article from the Industrial Track of Ada–Europe 2015.

Anonymous:2015:PWC

[860] Anonymous. Proceedings of the “workshop on challenges and new approaches for dependable and cyber-physical system engineering” of Ada–Europe 2015. Ada User Journal, 36(4):243, December 2015. CODEN AUJOET. ISSN 1381-6551.

Cancila:2015:E

[861] Daniela Cancila and Charles Robinson.

Editorial. Ada User Journal, 36(4): 244–245, December 2015. CODEN AU-JOET. ISSN 1381-6551.

Quadri:2015:MMC

[862] Imran Quadri, Alessandra Bagnato, Etienne Brosse, and Andrey Sadovykh. Modeling methodologies for cyberphysical systems: Research field study on inherent and future challenges. Ada User Journal, 36(4):246–253, December 2015. CODEN AUJOET. ISSN 1381-6551.

Pawlak:2015:DMC

[863] Adam Pawlak, Janusz Jezewski, and Krzysztof Horoba. Dependable medical cyber-physical system for home telecare of high-risk pregnancy. Ada User Journal, 36(4):254–258, December 2015. CO-DEN AUJOET. ISSN 1381-6551.

Burgio:2015:SNG

[864] Paolo Burgio, Carlos Alvarez, Eduard Ayguadé, Antonio Filgueras, Daniel Jiménez-González, Xavier Martorell, Nacho Navarro, and Roberto Giorgi. Simulating next-generation cyber-physical computing platforms. Ada User Journal, 36(4):259–263, December 2015. CODEN AUJOET. ISSN 1381-6551.

Mazzini:2015:CPO

[865] Silvia Mazzini. The CONCERTO Project: an open source methodology for designing, deploying, and operating reliable and safe CPS systems. Ada User Journal, 36(4):264–267, December 2015. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2015:AEAd

[866] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 36(4):268, December 2015. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2015:AESd

[867] Anonymous. Ada–Europe sponsors. Ada User Journal, 36(4):ibc, December 2015. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2016:EPA

[868] Anonymous. Editorial policy for Ada User Journal. Ada User Journal, 37(1): 2, March 2016. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2016:E

[869] Anonymous. Editorial. Ada User Journal, 37(1):3, March 2016. CODEN AU-JOET. ISSN 1381-6551.

Anonymous:2016:QND

[870] Anonymous. Quarterly news digest. Ada User Journal, 37(1):4–21, March 2016. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2016:CC

[871] Anonymous. Conference calendar. Ada User Journal, 37(1):22–27, March 2016. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2016:FE

[872] Anonymous. Forthcoming events. Ada User Journal, 37(1):28–30, March 2016. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2016:PRA

[873] Anonymous. Press release: Ada 2012 language standard corrigendum approved by ISO. Ada User Journal, 37(1): 31–34, March 2016. CODEN AUJOET. ISSN 1381-6551.

Barnes:2016:UA

[874] J. G. P. Barnes. Update for Ada 2012.

Ada User Journal, 37(1):35–50, March 2016. CODEN AUJOET. ISSN 1381-6551. Special Contribution.

Anonymous:2016:OIR

[875] Anonymous. Overview of the 17th International Real-Time Ada Workshop (IRTAW 2015). Ada User Journal, 37 (1):51, March 2016. CODEN AUJOET. ISSN 1381-6551.

Burns:2016:SSC

[876] A. Burns and J. A. de la Puente. Session summary: Conformance issues. Ada User Journal, 37(1):52, March 2016. CODEN AUJOET. ISSN 1381-6551.

Pinho:2016:SSF

[877] L. M. Pinho, S. Michell, and B. Moore.

Session summary: Fine-grained parallelism. Ada User Journal, 37(1):53–56, March 2016. CODEN AUJOET. ISSN 1381-6551.

Wellings:2016:SSL

[878] A. Wellings and J. Real. Session summary: Language abstractions. Ada User Journal, 37(1):57–58, March 2016. CO-DEN AUJOET. ISSN 1381-6551.

Burgio:2016:SNG

[879] Paolo Burgio, Carlos Alvarez, Eduard Ayguadé, Antonio Filgueras, Daniel Jiménez-González, Xavier Martorell, Nacho Navarro, and Roberto Giorgi. Simulating next-generation cyber-physical computing platforms. Ada User Journal, 37(1):59–63, March 2016. CODEN AUJOET. ISSN 1381-6551. Reprint of [864].

Anonymous:2016:AEA

[880] Anonymous. Ada–Europe associate members (national Ada organizations). Ada User Journal, 37(1):64–??, March 2016. CODEN AUJOET. ISSN 1381-6551.

Anonymous:2016:AES

[881] Anonymous. Ada–Europe sponsors. Ada User Journal, 37(1):ibc, March 2016. CODEN AUJOET. ISSN 1381-6551.